### v1

# Toroidal Cartesian Products Where One Factor is $3$-Connected

Abstract.
In this paper, we provide a partial classification of Cartesian products of graphs that embed on the torus where one factor is $3$-connected. We show that if $G\boxempty H$ embeds on the torus and $G$ is $3$-connected, then $H$ must be a path on two vertices, $P_{2}$, and we provide significant evidence that $G$ must be outer-cylindrical. At the same time, we show that if $G$ is outer-cylindrical, then $G\boxempty P_{2}$ embeds on the torus. This all leads us to conjecture that if $G$ is $3$-connected then $G\boxempty H$ embeds on the torus if and only if $G$ is outer-cylindrical and $H=P_{2}$. We also determine the genera of $K_{4}\boxempty P_{3}$ and $H_{i}\boxempty P_{2}$ where $H_{i}$ is a planar, $3$-connected, non-outer-cylindrical graph derived from a wheel graph $W_{i}$ via two vertex splits and one edge addition.
1. Introduction
Classifying embeddings of graphs on surfaces is a well-studied topic in topological graph theory. Here, we are interested in embeddings of a Cartesian product of two graphs $G$ and $H$, denoted $G\boxempty H$, which is defined to have vertex set $V(G)\times V(H)$ and a pair of vertices $(u,v),(u^{\prime},v^{\prime})\in V(G)\times V(H)$ determine an edge in $G\boxempty H$ if and only $u=u^{\prime}$ and $vv)\in E(H)$ or $uu^{\prime}\in E(G)$ and $v=v^{\prime}$. We refer the reader to [7] for further background on graph Cartesian products. The fiber structures of graph Cartesian products can sometimes be leveraged to efficiently build embeddings and determine genera of such graphs; see [10], [11], [13], [14], [15], [17], [20], [21] for some examples. In particular, it would be interesting to know if one can determine the set of all Cartesian products of graphs that embed on a fixed surface. Such a classification was established for planar graphs, that is, graphs that embed in the plane, in the work of Behzad and Mahmodian [4]. More recently, the second author along with Abell and McDermott have provided a complete classification of Cartesian products of graphs that embed in the projective plane in [1].
In this paper, we are interested in classifying which Cartesian product of graphs are toroidal, i.e., embed on the torus. While the classifications of planar and projective planar Cartesian products were reasonably straightforward, we expect the classification of toroidal Cartesian products to be far more complicated. To start, the work of Kuratowski [8] and Wagner [19] showed there are only two minor-minimal obstructions to planarity, namely $K_{5}$ and $K_{3,3}$, and Archdeacon in [3] showed that there is a complete list of 35 graphs that are minor-minimal obstructions to embedding on the projective plane. However, the list of minor-minimal obstructions for embedding on the torus isn’t completely known, and this list contains at least 17 thousand forbidden minors; see [12]. In addition, when restricting to Cartesian products of graphs, we see far more flexibility in toroidal embeddings than planar or projective planar embeddings with regards to (vertex) connectivity. If $G\boxempty H$ is planar or projective planar, then both $G$ and $H$ are each at most $2$-connected. This follows from the fact that $K_{4}\boxempty P_{2}$ is both non-planar and non-projective planar, and any $3$-connected graph contains a $K_{4}$ minor; see Lemma 4.2. However, in Section 5, we provide infinite classes of $3$-connected and $4$-connected graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal. Thus, even a partial classification of toroidal Cartesian products could be a far more complicated task than the planar and projective planar cases.
Here, we provide some major first steps towards this classification problem, with the additional assumption that one factor is $3$-connected. A simple combinatorial argument shows that if $G\boxempty H$ is toroidal, then $G$ and $H$ can each be at most $4$-connected in Proposition 4.1. Then in Theorem 4.3 we show that if $G$ is $3$-connected, then any toroidal $G\boxempty H$ must have $H=P_{2}$. This proof follows from the fact that $K_{4}\boxempty P_{3}$ is not toroidal, which is proved in Theorem 3.4. In Section 5, we provide two explicit constructions of infinite classes of toroidal $G_{i}\boxempty P_{2}$ and observe that any such $G_{i}$ in these constructions is planar and outer-cylindrical; see the beginning of Section 5 for the definition of outer-cylindrical. This all leads us to Conjecture 5.4, which states that if $G$ is $3$-connected, then $G\boxempty H$ is toroidal if and only if $H=P_{2}$ and and $G$ outer-cylindrical. Our work proves the backwards direction of this conjecture and provides strong evidence in favor of the forward direction. Theorem 4.3 already shows that $H=P_{2}$ holds under the hypotheses of the forward direction. Section 6 provides further evidence that the forward direction of this conjecture holds by showing the existence of an infinite set of planar, $3$-connected, non-outer-cylindrical graphs $\{H_{i}\}$ such that each $H_{i}\boxempty P_{2}$ does not embed on the torus and $H_{i}$ is “not far” in some sense from a wheel graph, $W_{i}$, which is an outer-cylindrical $3$-connected graph.
There are two significant by-products of our work on this classification problem. In Section 3 we prove that the genus of $K_{4}\boxempty P_{3}$ is two, highlighted in Theorem 3.4. This work requires a careful combinatorial analysis of all potential embeddings of this graph on a torus, which employs techniques similar to [5]. In particular, even though the girth of $K_{4}\boxempty P_{3}$ is three, one cannot hope for a triangular embedding since every edge in a $P_{3}$ fiber does not lie on a $3$-cycle. Thus, determining the genera of such a graph is not a straightforward task. In addition, in Section 6, we prove Theorem 6.10, which shows that $H\boxempty P_{2}$ has genus two, where $H$ is a specified planar, $3$-connected, non-outer-cylindrical graph. Parts of this proof follow the structure of the proof of Theorem 3.4, but the analysis is much more complex. As far as the authors can tell, both Theorem 3.4 and Theorem 6.10 provide novel contributions to the topological graph theory literature.
We would like to acknowledge the Furman University Mathematics Department for financially supporting this work via the Summer Mathematics Undergraduate Research Fellowships.
2. Background
In this paper, we assume all graphs are simple, connected, and non-trivial (in the sense of being nonempty and not a single vertex), unless otherwise noted. We use $S_{g}$ to denote an orientable, closed, connected surface of genus $g\in\mathbb{N}\cup\{0\}$. Given an embedding of a graph $G$ on a surface $S_{g}$ via a function $f:|G|\rightarrow S_{g}$ where $|G|$ represents a geometric realization of $G$, we sometimes use $f(G)$ to denote this embedding or just $G$ when the embedding is clear from context. If each component of $S_{g}\setminus f(G)$ is homeomorphic to an open disk, then we say this embedding is a $2$-cell embedding. The genus of a graph $G$, denoted $\gamma(G)$, is the minimum $g\in\mathbb{N}\cup\{0\}$ such that $G$ embeds on $S_{g}$. An embedding of $G$ that realizes the minimal genus is called a minimal embedding. We will mainly be interested in $2$-cell embeddings of graphs, and since the work of Youngs [22] shows that each minimal embedding of $G$ is a $2$-cell embedding, we usually don’t need to make a distinction when determining the genera of a graph. We refer the reader to [6] for further background on graph embeddings on surfaces.
A $2$-cell embedding $f:|G|\rightarrow S_{g}$ provides a $2$-cell decomposition of $S_{g}$ with $|V(G)|$ vertices, $|E(G)|$ edges, and $|F|$ faces, where each face is a component of $S_{g}\setminus f(G)$. Given such an embedding, the Euler characteristic of $S_{g}$ is given by $2-2g=\chi(S_{g})=|V(G)|-|E(G)|+|F|$. The following proposition is a well known lower bound on genus, coming from this Euler characteristic formula and analyzing the minimal number of edges on any face in an embedding of such a graph. Recall that the girth of a graph $G$ (that is not a tree) is the length of its shortest cycle.
Proposition 2.1.
Let $G$ be a graph (that is not a tree) with girth $\alpha$. Then
$\gamma(G)\geq 1+\frac{|E(G)|(\alpha-2)}{2\alpha}-\frac{|V(G)|}{2}.$
Equality holds if and only if there exists an embedding of $G$ where every face is bound by an $\alpha$-cycle and every edge borders two distinct faces.
Suppose we are given a $2$-cell embedding of $G$. We let $f_{i}$ denote the number of $i$-gon faces for this embedding. We use the terms triangle, quadrilateral, pentagon, and hexagon to refer to faces of a given embedding that are bound by a $3$-cycle, 44-cycle, 55-cycle, and $6$-cycle, respectively. We say that an embedding of a graph is a triangular embedding if every face is a triangle. We note the following basic but important facts about $2$-cell embeddings of $G$:
Fact 1: Each vertex of $G$ has an open $2$-cell neighborhood in $S$.
Fact 2: Each edge of $G$ borders precisely two (not necessarily unique) faces.
Fact 3: If $G=H\boxempty P_{n}$ for some graph $H$, then each edge from a $P_{n}$ fiber borders two $k$-gon faces, where $k\geq 4$.
Fact 4: $2|E(G)|=\displaystyle\sum_{i=3}^{\infty}if_{i}$.
Facts (1) and (2) are general facts about $2$-cell embeddings. Fact (3) follows from (2) and since any edge in a $P_{n}$ fiber is not contained in a $3$-cycle. Fact (4) follows from (2). This collection of facts will play an important role in ruling out potential toroidal embeddings of graphs in Section 3 and Section 6.
A graph $G^{\prime}$ is a minor of $G$ if $G^{\prime}$ can be obtained by a series of vertex deletions, edge deletions, or edge contractions on $G$. Another way to find a lower bound for the genus of a graph $G$ is by finding the genus (or lower bound on the genus) for a subgraph or minor $G$. Since we will be interested in Cartesian products of graphs in this paper, we first confirm that the product of minors from each factor results in a minor of the product.
Proposition 2.2.
If $G^{\prime}$ and $H^{\prime}$ are minors of $G$ and $H$ then $G^{\prime}\boxempty H^{\prime}$ is a minor of $G\boxempty H$.
Proof.
Let $G$ and $H$ be graphs and suppose $G^{\prime}$ be a minor of $G$. Recall that the vertex sets and edge sets of $G\boxempty H$ are defined as $V(G\boxempty H)=V(G)\times V(H)$ and $E(G\boxempty H)=(V(G)\times E(H))\cup(E(G)\times V(H))$. It suffices to show $G^{\prime}\boxempty H$ is a minor of $G\boxempty H$ under an edge deletion, vertex deletion, or edge contraction, which we now handle as three separate cases.
Case 1: Let $G^{\prime}=G-e$ for some $e\in E(G)$. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by deleting edge $(e,v)\in E(G)\times V(H)$, for each $v\in V(H)$.
Case 2: Let $G^{\prime}=G-v_{0}$ for some $v_{0}\in V(G)$. Recall that deleting some vertex $v_{0}$ also deletes all edges incident to $v_{0}$. Then we can obtain $G^{\prime}\boxempty H$ from $G\boxempty H$ by deleting vertex $(v_{0},v)\in V(G)\times V(H)$, for each $v\in V(H)$ (and deleting all its incident edges).
Case 3: Given edge $e\in E(G)$ with endpoints $a$ and $b$, we edge contract $e$ in $G$ to create $G^{\prime}$ which now has a new vertex $u\in V(G^{\prime})$ to which $a$ and $b$ have been identified. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by contracting edge $(e,v)\in E(G)\times V(H)$, for each $v\in V(H)$. ∎
Corollary 2.3.
Suppose $G^{\prime}$ and $H^{\prime}$ are minors (or subgraphs) of graphs $G$ and $H$, respectively. Then $\gamma(G\boxempty H)\geq\gamma(G^{\prime}\boxempty H^{\prime})$.
3. The genus of $K_{4}\boxempty P_{3}$
In this section, we will show that $\gamma(K_{4}\boxempty P_{3})=2$. Since every $3$-connected graph contains a $K_{4}$ minor (see Lemma 4.2), this will play a pivotal role in classifying toroidal Cartesian products where one of the factors is $3$-connected.
For the remainder of this section, set $G=K_{4}\boxempty P_{3}$. We have that $|V(G)|=12$ and $|E(G)|=26$. In addition, there are twelve $3$-cycles in $G$, with four $3$-cycles in each $K_{4}$ fiber. In particular, no edge contained in a $P_{3}$ fiber is contained in a $3$-cycle.
Lemma 3.1.
For any $2$-cell embedding of $G$, at most two $3$-cycles in each $K_{4}$ fiber are triangles, and so, $f_{3}\leq 6$.
Proof.
First, since $G$ has a girth of three, any triangle in a $2$-cell embedding of $G$ must be bound by a $3$-cycle. Suppose we are given a $2$-cell embedding of $G$ where there exist three triangles whose boundary $3$-cycles are all contained in the same $K_{4}$ fiber . Then there exists some vertex $v$ in this $K_{4}$ fiber where all three edges ($e_{1}$, $e_{2}$, and $e_{3}$) incident to $v$ in this $K_{4}$ fiber already border two faces in this embedding. Consider the unique edge $e\in E(K_{4}\boxempty P_{2})$ that is incident to $v$ and contained in a $P_{3}$ fiber. By Fact (2), $e$ must border exactly two faces in this embedding. However, any facial walk for such an embedding of $G$ that contains $e$ also contains one of the $e_{i}$ for $i=1,2,3$; otherwise, a vertex incident to $e$ would be degree $1$, which is not the case. However, this implies that some $e_{i}$ borders three faces, and so, violates Fact (1). Thus, any $2$-cell embedding of $G$ admits at most two triangles coming from any $K_{4}$ fiber. Since $3$-cycles of $G$ only occur within each $K_{4}$ fiber, we have that $f_{3}\leq 6$, as needed. ∎
This lemma already places strong restrictions on potential embeddings of $G$. Suppose $G$ has a $2$-cell embedding on $S_{g}$, for some $g\in\mathbb{N}\cup\{0\}$. Then $2-2g=\chi(S_{g})=12-26+|F|$, and so, $g=8-\frac{|F|}{2}$. From Fact 4, we know that given any $2$-cell embedding of $G$, we must have $2|E(G)|=52=\displaystyle\sum_{i=3}^{\infty}if_{i}$. In order to maximize the number of faces in order to minimize genus, one would attempt to use six triangles in an embedding, and so, $34=\displaystyle\sum_{i=4}^{\infty}if_{i}$. This implies that any $2$-cell embedding of $G$ has at most $14$ faces. Thus $\gamma(G)\geq 1$. Our main goal is to now prove that $\gamma(G)\neq 1$ by showing that an embedding of $G$ with $|F|=14$ is not possible.
Proposition 3.2.
Suppose $G$ admits a $2$-cell embedding on the torus. Then this embedding has $|F|=14$ with the following possibilities:
  1. (1)
    $f_{3}=4$ and $f_{4}=10$,
  2. (2)
    $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$,
  3. (3)
    $f_{3}=6$, $f_{4}=6$, and $f_{5}=2$, or
  4. (4)
    $f_{3}=6$, $f_{4}=7$, and $f_{6}=1$.
Proof.
As noted above, any toroidal embedding of $G$ must have $|F|=14$. By Fact (4), we have that
$52=\sum_{i=3}^{\infty}if_{i}=3f_{3}+\sum_{i=4}^{\infty}if_{i}\geq 3f_{3}+4\sum_{i=4}^{\infty}f_{i}.$
Thus,
$\frac{52-3f_{3}}{4}\geq\sum_{i=4}^{\infty}f_{i}=|F|-f_{3}.$
This implies $13+\frac{f_{3}}{4}\geq|F|$. Since $|F|=14$, this implies that $f_{3}\geq 4$. We will frequently make use of this type of combinatorial restriction in this proof, though we will suppress most of the computations moving forward. Combined with Lemma 3.1, we have $4\leq f_{3}\leq 6$.
First, suppose $f_{3}=6$. Then $\sum_{i=4}^{\infty}f_{i}=8$ and $\sum_{i=4}^{\infty}if_{i}=34$. As a result, at least six quadrilaterals must be used. If $f_{4}=6$, then the remaining two faces must be pentagons. If $f_{4}=7$, then the remaining face must be a hexagon. If $f_{4}=8$ then we would have $\sum_{i=4}^{\infty}if_{i}=32$, which is a contradiction.
Now suppose $f_{3}=5$. Then $\sum_{i=4}^{\infty}f_{i}=9$ and $\sum_{i=4}^{\infty}if_{i}=37$. This forces eight quadrilaterals to be used. Thus, $f_{4}=8$ and $f_{5}=1$.
Finally, suppose $f_{3}=4$. Then $\sum_{i=4}^{\infty}f_{i}=10$ and $\sum_{i=4}^{\infty}if_{i}=40$, which can only occur if all ten faces are quadrilaterals. ∎
In light of Proposition 3.2, it will help to more carefully describe how faces can behave in a $2$-cell embedding of $G$ on the torus. We first introduce some necessary notation. We say that a $K_{4}$ fiber of $G$ is an outer $K_{4}$ if its vertices are adjacent to exactly four of the edges contained in the $P_{3}$ fibers. The unique $K_{4}$ fiber of $G$ whose vertices are adjacent to all edges in the $P_{3}$ fibers is called the inner $K_{4}$ fiber. Label the $K_{4}$ fibers $K^{1}$, $K^{2}$, and $K^{3}$, respectively, with $K^{2}$ inner. We use $E_{1}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{1}$ and $K^{2}$ and $E_{2}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{2}$ and $K^{3}$. Given a $2$-cell embedding of $G$, we say that a quadrilateral is a mixed quadrilateral if it’s boundary is a $4$-cycle of $G$ that consists of two edges from $E_{1}$ ($E_{2}$), one edge from a $K^{1}$ ($K^{3}$), and the corresponding edge from $K^{2}$. Note, that any quadrilateral in an embedding of $G$ that contains a $P_{3}$ fiber edge must be a mixed quadrilateral. Up to symmetry, there are two distinct types of pentagons that bound $5$-cycles in $G$ that could occur in this $2$-cell embedding of $G$, which we denote $p_{1}$ and $p_{2}$. The 5-cycle boundary of $p_{1}$ contains one edge on $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and two edges in an outer $K_{4}$. Similarly the 5-cycle boundary of $p_{2}$ contains two edges on $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and one edge in an outer $K_{4}$. Hexagons that bound $6$-cycles in $G$ could possibly be placed in four distinct ways in a $2$-cell embedding of $G$, denoted $h_{1}$, $h_{2}$, $h_{3}$, and $h_{4}$. The 6-cycle boundary of $h_{1}$ contains one edge in $K^{1}$, two edges in $E_{1}$, two edges in $E_{2}$, and one edge in $K^{3}$. The 6-cycle boundary of $h_{2}$ contains two edges in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and two edges in an outer $K_{4}$. The 6-cycle boundary of $h_{3}$ contains one edge in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K_{4}$. The $6$-cycle boundary of $h_{4}$ contains one edge in an outer $K_{4}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K^{2}$.
For the above description, we only considered faces in a $2$-cell embedding of $G$ on the torus whose boundary was an appropriate cycle. We now show those are the only types of faces that can occur in such an embedding.
Lemma 3.3.
If $G$ has a $2$-cell embedding on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle of $G$.
Proof.
Suppose we are given a $2$-cell embedding of $G$ on the torus. Let $F$ be a face in this embedding. Then the boundary of $F$ is a closed walk in $G$ determined by a “facial walk” of this embedding. From Proposition 3.2, we only need to consider when $F$ is a triangle, quadrilateral, pentagon, or hexagon. We will show that if the facial walk for $F$ isn’t an appropriate cycle, then we always reach a contradiction.
First, suppose our embedding admits an $n$-gon face, $F$, that is not bound by an $n$-cycle for some $n=3$, 44, or $5$. The boundary of $F$ is must be a closed walk in $G$, and since the girth of $G$ is three, this closed walk must have a subwalk of the form eveeve for some $e\in E(G)$ and $v\in V(G)$. However, this implies that $G$ admits a degree one vertex, which is a contradiction.
Now, suppose $F$ is a hexagon and the facial walk of $F$ determines a closed walk of length six in $G$ that isn’t a $6$-cycle. If the facial walk for $F$ doesn’t contain any cycle subgraphs, then we once again get a degree one vertex in $G$, which is a contradiction. However, in any $K_{4}$ fiber of $G$, there exist pairs of $3$-cycles that intersect along a single edge. This is the only scenario where the facial walk of $F$ could contain cycle subgraphs and not be a $6$-cycle. However, such an $F$ is not possible. If $f_{6}\geq 1$, then Proposition 3.2 shows that $f_{3}=6$ and $f_{4}=7$. Note that, any triangle in this embedding can only bound edges in $K_{4}$ fibers, and we just assumed our one hexagon only bounds edges in $K_{4}$ fibers. There are only seven quadrilaterals left in this embedding and they must be mixed quadrilaterals, each of which meets exactly two $P_{3}$ fiber edges. However, there are eight edges in our $P_{3}$ fibers, each of which must lie on the boundary of two faces in this embedding, and so, some edge in our $P_{3}$ fiber will only be border one face, which is a contradiction. ∎
Theorem 3.4.
$\gamma(K_{4}\boxempty P_{3})=2$.
Proof.
We will first show that $G=K_{4}\boxempty P_{3}$ does not embed on the torus by analyzing the four possibilities given in Proposition 3.2. We then show that $G$ does admit an embedding on $S_{2}$.
For sake of contradiction, suppose $G$ 22-cell embeds on the torus with $|F|=14$. In each case, we show that our assumpotions violate Fact (2), giving a necessary contradiction. The following is an important fact we will use throughout the proof: any $e\in E_{1}\cup E_{2}$ does not lie on a $3$-cycle and if such an edge $e$ is bound by a quadrilateral in this embedding, then it must be a mixed quadrilateral.
Suppose $f_{3}=4$ and $f_{4}=10$. To make sure each edge in $E_{1}\cup E_{2}$ is incident to two faces, this embedding must have eight mixed quadrilaterals, four of which are incident to edges in $E_{1}$ and four of which are incident to edges in $E_{2}$. Each of these eight mixed quadrilaterals is incident to exactly one edge in $K^{2}$. Since $12=2|E(K^{2})|$ and each edge must border exactly two faces, four of the (not necessarily distinct) edges in $K^{2}$ must be bound by faces in this embedding that are not mixed quadrilaterals. Since each edge in $E_{1}\cup E_{2}$ already bounds two faces, there must exist a $4$-cycle in $K^{2}$ that is the boundary of a quadrilateral in this embedding. At this point, all faces incident to edges in $K^{2}$ have been determined. Thus, by Lemma 3.1 and our assumptions on $f_{3}$ and $f_{4}$, there are two triangles incident to only edges in $K^{1}$, two triangles incident to only edges in $K^{3}$, and one quadrilateral that must be incident to either only edges in $K^{1}$ or $K^{3}$. Note that $K^{1}$ and $K^{3}$ each have four edges incident to mixed quadrilaterals. Thus, either some edges in $K^{1}$ don’t border two faces and some edges in $K^{2}$ border at least three faces, or vice versa, giving a contradiction.
Suppose $|F|=14$ and $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$. The two types of pentagons ($p_{1}$ and $p_{2}$ described earlier in this section) each bound two edges coming from either $E_{1}$ or $E_{2}$. As a result, we must then have seven mixed quadrilaterals in this embedding to make sure each edge in $E_{1}\cup E_{2}$ borders exactly two faces. This leaves one quadrilateral and five triangles to complete the embedding. Since each mixed quadrilateral is incident to an edge in $K^{2}$ and a pentagon is incident to either one or two edges in $K^{2}$, either four or three more (not necessarily distinct) edges in $K^{2}$ must be incident to the other faces in this embedding. By Lemma 3.1 and $f_{3}=5$, either one or two triangles are incident to only edges in $K^{2}$. Having two triangles contained in $K^{2}$ results in some edge in $K^{2}$ having at least three faces incident to it, which is a contradiction. Having only one triangle contained in $K^{2}$ results in some edge in $K^{2}$ having only one face incident to it, which is a contradiction. The only other possibility is that one triangle and the one (non-mixed) quadrilateral are incident to edges in $K^{2}$, but this would again result in some edge of $K^{2}$ having three faces incident to it. Thus, an embedding of $G$ with $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$ is not possible.
Suppose $f_{3}=6$, $f_{4}=6$, and $f_{5}=2$. By Lemma 3.1, each $K_{4}$ fiber contains exactly two triangles. Note that, each pentagon borders a pair of edges in either $E_{1}$ or $E_{2}$. Thus, all six quadrilaterals must be mixed quadrilaterals in order to make sure each edge in $E_{1}\cup E_{2}$ borders exactly two faces. Each mixed quadrilateral borders exactly one edge in $K^{2}$, each pentagon borders either one or two edges in $K^{2}$, and there are two triangles bordering a total of six edges in $K^{2}$. As a result, a minimum of $14$ edges in $K^{2}$ are bordered by faces, which forces some edge in $K^{2}$ to border at least three faces, providing a contradiction.
Suppose $f_{3}=6$, $f_{4}=7$, and $f_{6}=1$. By Lemma 3.1, each $K_{4}$ fiber contains exactly two triangles. As discussed before the proof, there are three types of hexagons that can exist in an embedding of $G$. We now break this case into sub-cases, depending on what type of hexagon is used.
First, suppose an $h_{1}$ hexagon is used in this embedding. Since this hexagon borders four edges from $E_{1}\cup E_{2}$, we must have six mixed quadrilaterals in this embedding to ensure every edge in $E_{1}\cup E_{2}$ borders exactly two faces. Thus, we only need to determine where the last quadrilateral is placed in this embedding. This quadrilateral must be bounded by edges in a $K_{4}$ fiber since the union of the pentagon and six mixed quadrilaterals collectively border every edge in $E_{1}\cup E_{2}$ twice. However, no matter which $K_{4}$ fiber is used here, we will have an edge incident to at least three faces, providing a contradiction.
Now, suppose an $h_{2}$, $h_{3}$, or $h_{4}$ hexagon is used in this embedding. Such a hexagon does not border one of the outer $K_{4}$ fibers, say $K^{3}$. In addition, such a hexagon borders exactly two edges from either $E_{1}$ or $E_{2}$. As a result, all seven quadrilaterals must all be mixed quadrilaterals to ensure every edge in $E_{1}\cup E_{2}$ borders exactly two faces. Four of these mixed quadrilaterals will each be incident to one edge in $K^{3}$ and the other three mixed quadrilaterals do not border $K^{3}$. At this point, we have determined where all our faces are placed and we see that some edges in $K^{3}$ are bordered by at most one face, which is a contradiction.
We now describe how to embed $G$ on $S_{2}$. First, embed each $K_{4}$ fiber on a separate $2$-sphere as the $1$-skeleton of a tetrahedron. We use $F_{i}$ for $i=1,\ldots,4$ to denote the four triangles in one of these embeddings. First, delete the interiors of $F_{1}$ and $F_{2}$ from the $K^{1}$ embedding and from the $K^{2}$ embedding. Insert a tube (homeomorphic to $\mathbb{S}^{1}\times[0,1]$), between the $F_{j}$ face deleted from $K^{1}$ and the $F_{j}$ face deleted from $K^{2}$ for $j=1,2$. The four edges of $E_{1}$ can be embedded on these two tubes. Now, delete the triangular faces $F_{3}$ and $F_{4}$ coming from $K^{2}$’s embedding on a $2$-sphere and also delete $F_{3}$ and $F_{4}$ from $K^{3}$’s embedding on a $2$-sphere. Insert two tubes in a similar fashion to embed the edges of $E_{2}$, completing the embedding of $G$. Since this embedding involves three disjoint $2$-spheres joined by four tubes, the resulting genus of the surface is two. ∎
4. Restricting toroidal $G\boxempty H$ via vertex connectivity.
The vertex-connectivity of a graph $G$, denoted $\kappa(G)$, is the minimum cardinality of a vertex-cut if $G$ is not complete or $\kappa(G)=n-1$ if $G=K_{n}$ for $n\in\mathbb{N}$. A connected graph $G$ is $k$-connected for $1\leq k\leq\kappa(G)$. Our goal in this section is to examine how vertex-connectivity can affect whether or not $G\boxempty H$ is toroidal. We first note that $G$ and $H$ can be at most $4$-connected, if $G\boxempty H$ is toroidal.
Proposition 4.1.
If $\gamma(G\boxempty H)\leq 1$, then $\kappa(G),\kappa(H)\leq 4$.
Proof.
We proceed by contrapositive. Suppose $\kappa(G)\geq 5$. Since $H$ is a simple connected graph with $|V(H)|\geq 2$, we see that $H$ contains a $P_{2}$ subgraph. Note that, $|V(G\boxempty P_{2})|=2|V(G)|$ and $|E(G\boxempty P_{2})|=2|E(G)|+|V(G)|$. Since $G$ is simple, the girth of $G\boxempty P_{2}$ is at least three. By Proposition 2.1, we have that
$\gamma(G\boxempty P_{2})\geq 1+\frac{2|E(G)|+|V(G)|}{6}-\frac{2|V(G)|}{2}=1+\frac{1}{3}|E(G)|-\frac{5}{6}|V(G)|.$
Since $G$ is $5$-connected, for all $v\in V(G)$, we have $deg(v)\geq 5$. Since $G$ is simple, every edge is incident to two distinct vertices. Thus, $\frac{5}{2}|V(G)|\leq|E(G)|$, and so, we can conclude that
$\gamma(G\boxempty P_{2})\geq 1+\frac{1}{3}(\frac{5}{2}|V(G)|)-\frac{5}{6}|V(G)|=1.$
Equality holds if and only if $G$ is $5$-regular and $G\boxempty P_{2}$ admits a triangular emdedding. However, any edge in a $P_{2}$ fiber is not contained in any $3$-cycle of $G\boxempty P_{2}$, and so, a triangular embedding of $G\boxempty P_{2}$ does not exist. Therefore, $\gamma(G\boxempty H)\geq\gamma(G\boxempty P_{2})>1.$
We now work to show that severe restrictions can be placed on $H$ when $\kappa(G)\in\{3,4\}$. We first need a basic fact about $3$-connected graphs.
Lemma 4.2.
Let $G$ be a $3$-connected graph. Then $G$ contains a $K_{4}$ minor.
Proof.
The work of Tutte [18] shows that any simple, $3$-connected graph contains a wheel graph $W_{n}$ as a minor, for some $n\geq 3$; see Figure 2 for a diagram of $W_{5}$. At the same time, $W_{m}$ is a minor of $W_{n}$, whenever $m\leq n$. It is easy to see that one can first delete an appropriate number of the spoke edges of $W_{n}$ that connect to the central vertex, and then, contract some of the edges along the $n$-cycle of $W_{n}$ to obtain a $W_{m}$ minor. Thus, $W_{3}=K_{4}$ is a minor of any simple, $3$-connected graph, as needed. ∎
This result has immediate consequences for classifying toroidal Cartesian products where one factor is $3$-connected.
Theorem 4.3.
Let $G$ be a $3$-connected graph. If $\gamma(G\boxempty H)\leq 1$, then $H=P_{2}$.
Proof.
We proceed by contrapositive. Suppose $H$ is a (simple, connected, non-trivial) graph where $H\neq P_{2}$. Then $H$ contains a $P_{3}$ subgraph. By Lemma 4.2, $G$ has a $K_{4}$ minor. Then by Theorem 3.4 and Corollary 2.3, we have that $\gamma(G\boxempty H)\geq\gamma(K_{4}\boxempty P_{3})=2$. ∎
Combined, Theorem 4.3 and Proposition 4.1 give rise to the following questions:
Question 4.4.
Does there exist an infinite set of $3$-connected (44-connected) graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal? If so, can we classify such $G_{i}$?
5. Constructing toroidal $G\boxempty P_{2}$, where $\kappa(G)\in\{3,4\}$.
In this section, we address the first part of Question 4.4 and provide a conjecture for the second part of Question 4.4. This first requires us to introduce a new class of graphs. A graph $G$ is outer-cylindrical if it embeds on the $2$-sphere such that there are two distinct faces whose boundaries collectively contain all the vertices of $G$. Since the class of outer-cylindrical graphs is closed under taking minors, the Graph Minor Theorem of Robertson and Seymour [16] implies that there is a finite list of minor-minimal obstructions for this class of graphs, which were then determined by Archdeacon–Bonnington–Dean–Hartsfield–Scott [2]. Following their notation, we will use the abbreviation OC to stand for outer-cylindrical and NOC to stand for non-outer-cylindrical. The following proposition motivates our interest in OC graphs.
Proposition 5.1.
Let $G$ be an OC-graph. Then $\gamma(G\boxempty P_{2})\leq 1$.
Proof.
We first construct disjoint OC-embeddings of the two $G$ fibers (call them $G_{1}$ and $G_{2}$) of $G\boxempty P_{2}$ in the $2$-sphere $\mathbb{R}\cup\{\infty\}$. Take an OC-embedding of $G$ as $G_{1}$, and reflect $G_{1}$ over a line that does not intersect $G_{1}$ to build the OC-embedding of $G_{2}$. All the vertices of each $G_{i}$ for $i=1,2$ are contained in two faces of this embedding, and one of these faces, which we call the outer face, is common between $G_{1}$ and $G_{2}$. For each vertex $v$ in $G_{1}$ on this outer face, we can embed on edge on the $2$-sphere connecting $v$ to the corresponding vertex in $G_{2}$. We can then cut out the interiors of the “inner” faces for $G_{1}$ and $G_{2}$ and add a tube between them carrying the remaining edges of $G\boxempty P_{2}$ to build an embedding of $G\boxempty P_{2}$ on the torus.∎
[…]
Figure 1. The inner and outer faces for the $G_{1}$ fiber are depicted in green and the inner and outer faces for the $G_{2}$ fiber are depicted in black. The edges contained in the $P_{2}$ fibers that connect to vertices on the two outer faces are depicted in blue. The edges contained in the $P_{2}$ fibers that connect to vertices on the inner faces are not depicted.
It is easy to see that there exist infinite classes of $3$-connected graphs that are OC. For this, we recall the definition of a wheel of order n for $n\geq 3$, denoted $W_{n}$, as the simple connected graph with $n+1$ vertices and $2n$ edges with the following properties:
  1. (1)
    $V(W_{n})=\{c,b_{0},\ldots,b_{n-1}\}$,
  2. (2)
    $c$ is adjacent to each $b_{i}$ for $i=0,\ldots n-1$,
  3. (3)
    $b_{i}$ and $b_{j}$ are adjacent if and only if $|i-j|=1\mod n$.
[…]
Figure 2. The standard planar embedding of $W_{5}$ with vertices labeled.
Proposition 5.2.
For $n\geq 3$, each wheel $W_{n}$ is OC. Consequently, there exists an infinite set of OC, $3$-connected graph $\{W_{n}\}_{n=3}^{\infty}$ such that each $W_{n}\boxempty P_{2}$ is toroidal.
Proof.
For $n\geq 3$, consider the standard planar embedding of $W_{n}$ as given in Figure 2. Then the outer region is a face bound by the $n$-cycle $b_{0}b_{1},\ldots,b_{n-1}$. The triangle $cb_{0}b_{1}$ bounds $c$, and so, these two faces contain every vertex, which implies that this embedding is OC. The work of Tutte [18] shows that each wheel is simple and $3$-connected. By Proposition 5.1, each $W_{n}\boxempty P_{2}$ is toroidal. ∎
We can also construct $4$-connected graphs that are OC. For $n\geq 4$, define $C^{2}_{n}$ as the graph obtained from the $n$-cycle $C_{n}$ by adding an edge between every pair of vertices that are a distance two apart in $C_{n}$. See Figure 3 for a standard planar embedding of $C^{2}_{6}$.
[…]
Figure 3. The standard planar embedding of $C^{2}_{6}$ so that the green cycle $v_{1}v_{3}v_{5}$ and red cycle $v_{2}v_{4}v_{6}$ each bound faces for an OC embedding.
Proposition 5.3.
For $n\geq 3$, each wheel $C^{2}_{2n}$ is OC. Consequently, there exists an infinite set of OC, $4$-connected graph $\{C^{2}_{2n}\}_{n=3}^{\infty}$ such that each $C^{2}_{2n}\boxempty P_{2}$ is toroidal.
Proof.
Provide a cyclic labeling of the vertices of cycle $C_{2n}$ as $v_{1},\ldots,v_{2n}$. Consider the standard planar embedding of the cycle $C_{2n}$, which partitions the plane into two regions, an inner region and an outer region. One can add in the edges of $C^{2}_{2n}$ that are a distance two apart between odd vertices on the inner region and add in the edges of $C^{2}_{2n}$ that are a distance two apart between even vertices on the outer region so that this gives a planar embedding of $C^{2}_{2n}$ where one face (contained inside the inner region) has boundary that contains all the odd vertices and one face (contained in the outer region) has boundary that contains all even vertices; see Figure 3. Thus, $C^{2}_{2n}$ is OC. For $n\geq 3$, $C^{2}_{2n}$ is known to be $4$-connected; see [9]. By Proposition 5.1, each $C^{2}_{2n}\boxempty P_{2}$ is toroidal for $n\geq 3$. ∎
The work accomplished over the last two sections leads us to the following conjecture.
Conjecture 5.4.
Suppose $G$ is $3$-connected. Then $\gamma(G\boxempty H)\leq 1$ if and only if $G$ is OC and $H=P_{2}$.
We have already proved the backwards direction of Conjecture 5.4. We also have shown that if $G$ is $3$-connected and $\gamma(G\boxempty H)\leq 1$, then $H=P_{2}$. Furthermore, if a graph $G$ is non-planar, then $\gamma(G\boxempty H)>1$ since $G\boxempty H$ contains two disjoint copies of $G$. This leads to the following question:
Question 5.5.
If $G$ is a planar, $3$-connected, NOC graph, then must $\gamma(G\boxempty P_{2})>1?$
In the following section, we provide evidence showing that we don’t expect such $G\boxempty P_{2}$ to be toroidal, further supporting Conjecture 5.4.
6. $G\boxempty P_{2}$ where $G$ is planar, $3$-connected, and NOC.
Here, we provide an explicit construction for an infinite set of planar, $3$-connected, NOC graphs $\{H_{n}\}$, and then show that $H_{n}\boxempty P_{2}$ is non-toroidal. The construction of $H_{5}$ from a sequence of two vertex splits and one edge addition from the wheel $W_{5}$ is given in Figure 4. This process can be generalized to construct $H_{n}$ from $W_{n}$, for $n\geq 5$. Moving forward, we refer to the operations of vertex splitting and edge addition collectively as Tutte moves since Tutte used these operations in [18] to build any $3$-connected graph from some wheel graph.
[…]
Figure 4. A sequence of three Tutte moves on $W_{5}$ to create a planar, 33-connected NOC graph, $H=H_{5}$.
Theorem 6.1.
For each $n\geq 5$, there exist a sequence of three Tutte moves on $W_{n}$ that results in a planar, 3-connected, NOC graph $H_{n}$.
Proof.
Consider $W_{n}$, with the vertex labeling for a wheel graph given in Section 5. Perform a vertex split on $c$, creating two new vertices $c_{1}$ and $c_{2}$ with an edge connecting them. Further, assume $c_{1}$ is degree $3$ and adjacent to $c_{2}$, $b_{0}$, and $b_{1}$, while $c_{2}$ is degree n1n-1 and adjacent to $c_{1}$ and $b_{2},\ldots,b_{n-1}$. Next, add an edge ($e$ depicted in red in in part 3 of Figure 4) from $c_{2}$ to $b_{1}$. Finally, split $c_{2}$ into $c_{3}$ and $c_{4}$, such that $c_{3}$ is degree $4$ and adjacent to $b_{1}$, $c_{1}$, $b_{n-1}$, and $c_{4}$, while $c_{4}$ is degree n2n-2 and adjacent to $c_{3}$ and $b_{2},\ldots,b_{n-2}$.
We now check that each $H_{n}$ with $n\geq 5$ has the desired properties. Since $W_{n}$ is a $3$-connected graph and Tutte moves preserve being $3$-connected, we also have that $H_{n}$ is a $3$-connected graph. In addition, one can start with the standard wheel diagram for $W_{n}$, which is planar, and perform the necessary Tutte moves in a way that preserves planarity to create a planar diagram of $H_{n}$, as shown in Figure 4. Next, we see that $H_{5}$ contains a minor that is a minor-minimal NOC graph, specifically $\gamma_{2}$ of Figure 12 in [2]. This shows $H_{5}$ is NOC. Furthermore, we see that $H_{5}$ is a minor of $H_{n}$ (with $n\geq 5$) since one can delete the edges connecting $c_{4}$ to $b_{4},\ldots,b_{n-2}$ from $H_{n}$ to obtain a graph homeomorphic to $H_{5}$. Thus, $\gamma_{2}$ is a minor of each such $H_{n}$ with $n\geq 5$. Since the class of OC graphs is closed under taking minors, this shows that $H_{n}$ is NOC, for each $n\geq 5$. ∎
For the remainder of Section 6, we will refer to the planar, 3-connected, NOC graph depicted in step 4 of Figure 4 as $H=H_{5}$ and $G=\ H\boxempty P_{2}$ to simplify notation. Since $H$ is a minor of all graphs obtained via the construction outlined in Theorem 6.1, the genus of $G$ provides a lower bound on the genus of each $H_{n}\boxempty P_{2}$, for $n\geq 5$. This motivates us to prove that $G$ is non-toroidal, which we will do by the end of this section in Theorem 6.10. First, we note an important corollary of this theorem.
Corollary 6.2.
$\{H_{n}\}_{n=5}^{\infty}$ is an infinite set of planar, $3$-connected, NOC graphs such that each $H_{n}\boxempty P_{2}$ is non-toroidal.
Our corollary also implies that any $3$-connected graph $H^{\prime}$ that contains an $H_{n}$ minor will result in $H^{\prime}\boxempty P_{2}$ being non-toroidal. Recall that Tutte [18] showed that every $3$-connected graph is constructed via a finite series of a Tutte moves from some wheel $W_{n}$ for $n\geq 4$. Since $H_{n}$ is constructed via just three Tutte moves on $W_{n}$, we see that many $3$-connected graphs will meet the necessary qualities to be a minor of $H$. Thus, this work provides further evidence in support of Conjecture 5.4.
Mirroring the technique used for Theorem 3.4, we now prove the following proposition and lemma that describe the behavior of faces of a $2$-cell embedding of $G$ on the torus.
Proposition 6.3.
Suppose $G$ admits a $2$-cell embedding on a torus. This embedding has $|F|=18$ with the following possibilities:
  1. (1)
    $f_{3}=6$, $f_{4}=11$, and $f_{6}=1$
  2. (2)
    $f_{3}=6$, $f_{4}=10$, and $f_{5}=2$
  3. (3)
    $f_{3}=5$, $f_{4}=12$, and $f_{5}=1$, or
  4. (4)
    $f_{3}=4$ and $f_{4}=14$.
Proof.
Note that, $|E(G)|=34$ and $|V(G)|=16$. Thus, if $G$ admits a $2$-cell embedding on a torus, then this embedding must have $|F|=18$.
We now consider the possibilities for triangular faces in this embedding. Since the girth of $G$ is three, any triangle in this embedding must be bound by a $3$-cycle of $G$. We know no edge in the $P_{2}$ fibers is contained in a $3$-cycle, and so, every triangular face in $G$ must bound a $3$-cycle in an $H$-fiber. This implies $f_{3}\leq 6$ since each $H$-fiber contains only three $3$-cycles. If $f_{3}\leq 3$, then we reach a contradiction because of $2|E(G)|=68=\displaystyle\sum_{i=3}^{\infty}if_{i}$ and $|F|=18$. Thus, $4\leq f_{3}\leq f_{6}$. From here, one can explicitly check the three possible cases for $f_{3}$ lead to face sets (1)-(4) given in the statement of this Proposition. The details are very similar to the proof of Proposition 3.2, which we leave to the reader. ∎
Lemma 6.4.
If $G$ 22-cell embeds on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle in $G$.
Proof.
This proof follows the same type of argument given in Lemma 6.4. We leave the details to the reader. ∎
To prove Theorem 6.10, we now introduce specific notation used to describe $G=H\boxempty P_{2}$ for the remainder of this section. We denote the two $H$-fibers of $G$ as $H^{+}$ and $H^{-}$. We say that edges $e^{+}\in E(H^{+})$ and $e^{-}\in E(H^{-})$ are corresponding if their images are identical under the projection map $\pi:H\boxempty P_{2}\rightarrow H$. Let $E_{p}$ be the set of all edges in the $P_{2}$ fibers of $G$. Let $Q$ be the set of all 4-cycles in $G$ such that if $c\in Q$, then $c$ contains two distinct edges in $E_{p}$, one edge in $H^{+}$, and the corresponding edge in $H^{-}$. In fact, given any $e\in E_{p}$, the only $4$-cycles that contain $e$ are elements of $Q$. Note that $Q$ is in bijective correspondence with the edges of $H$. Given an embedding of $G$, we say that faces bound by cycles in $Q$ are mixed quadrilaterals. Suppose $G$ 2-cell embeds on some surface $S_{g}$ where for any edge $e\in E_{p}$, we see that $e$ is incident to two (mixed) quadrilateral faces. Then we let $Q^{*}\subseteq Q$ represent the set of 4-cycles that collectively bound any mixed quadrilaterals in this embedding. We set $E^{+}=Q^{*}\cap H^{+}$, and likewise, $E^{-}=Q^{*}\cap H^{-}$.
Lemma 6.5.
Suppose $G$ 2-cell embeds on some surface $S_{g}$ where for any edge $\in E_{p}, ee$ is incident to two (mixed) quadrilateral faces. Then the union of elements in $E^{+}$ form a disjoint union of cycles in $H^{+}$ that collectively visit every vertex in $H^{+}$. Similarly, the union of elements in $E^{-}$ form a disjoint union of cycles in $H^{-}$ that collectively visit every vertex in $H^{-}$.
Proof.
Let $v$ be some vertex in $H^{+}$. Then $v$ is incident to a unique edge in $E_{p}$ which is contained in exactly two distinct $4$-cycles in $Q^{*}$. By definition, each $Q^{*}$ cycle contains a distinct edge in $E^{+}$ also incident to $v$. Thus, every vertex in $H^{+}$ is incident to exactly two distinct edges in $E^{+}$, meaning $E^{+}$ creates a collection of one or more disjoint cycles whose union intersects every vertex in $H^{+}$. A similar argument can be replicated for $E^{-}$. ∎
We will consider two possible cases for a $2$-cell embedding of $G$ on the torus. In the first case, we will assume every edge in $E_{p}$ lies on the boundary of a (mixed) quadrilateral face in this embedding. By Proposition 6.3 and the fact that no edge in $E_{p}$ lies on a $3$-cycle of $G$, our second case only needs to consider when only a few edges in $E_{p}$ lies on the boundary of pentagon or hexagon and the rest of the edges of $E_{p}$ lie on the boundary of mixed quadrilaterals. We refer to such pentagons as mixed pentagons, and similarly we refer to such hexagons as mixed hexagons. Collectively, we refer to mixed quadrilaterals, pentagons, and hexagons of a $2$-cell embedding of $G$ as mixed $n$-gons.
For now, if we assume every edge in $E_{p}$ lies on the boundary of a (mixed) quadrilateral, then Lemma 6.5 shows the union of elements in $^{+}/EE^{-}$ form a collection of pairwise disjoint cycles $\{C_{i}\}$ in $H$ such that $\sum i=8$. This allows us to provide more specification to potential toroidal embeddings of $G$ in terms of such cycle collections. Note that, $H$ is a simple graph and one can check that $H$ does not contain a pair of disjoint $4$-cycles. Thus, any such collection $\{C_{i}\}$ contains either one $8$-cycle or a $5$-cycle and a $3$-cycle. Below, we list these possibilities following the vertex labeling given in Figure 5, and refer to these as cycle sets (of $E^{\pm}$).
8-cycles in $H$:
  1. (a)
    ($v_{1}$, $v_{2}$, $v_{3}$, $v_{6}$, $v_{7}$, $v_{4}$, $v_{8}$, $v_{5}$)
  2. (b)
    ($v_{1}$, $v_{2}$, $v_{3}$, $v_{6}$, $v_{7}$, $v_{8}$, $v_{4}$, $v_{5}$)
  3. (c)
    ($v_{1}$, $v_{2}$, $v_{6}$, $v_{3}$, $v_{4}$, $v_{7}$, $v_{8}$, $v_{5}$)
  4. (d)
    ($v_{1}$, $v_{7}$, $v_{6}$, $v_{2}$, $v_{3}$, $v_{4}$, $v_{8}$, $v_{5}$)
  5. (e)
    ($v_{1}$, $v_{2}$, $v_{6}$, $v_{3}$, $v_{4}$, $v_{5}$, $v_{8}$, $v_{7}$)
Pairs of disjoint 3- and 5-cycles in $H$:
  1. (f)
    ($v_{2}$, $v_{3}$, $v_{6}$), ($v_{1}$, $v_{7}$, $v_{4}$, $v_{8}$, $v_{5}$)
  2. (g)
    ($v_{4}$, $v_{5}$, $v_{8}$), ($v_{1}$, $v_{2}$, $v_{3}$, $v_{6}$, $v_{7}$)
[…]
Figure 5. The graph $H$ with vertex labels.
Lemma 6.6.
Suppose $G$ 22-cell embeds on the torus where both copies of the edge $(v_{4},v_{8})$ in the $H$ fibers each lie on the boundary of some mixed $n$-gon face (not necessarily the same one). Then $f_{3}=4$ for this embedding and in both of the $H$ fibers, the cycle $(v_{4},v_{7},v_{8})$ or the cycle $(v_{4},v_{5},v_{8})$ can border a triangular face, but not both.
Proof.
For sake of contradiction, suppose there exists an embedding of $G$ on the torus with at least five triangles, and both copies of the edge $(v_{4},v_{8})$ in the $H$ fibers each lie on the boundary of some mixed $n$-gon face (not necessarily the same one). Note, $G$ contains exactly six $3$-cycles, three in $H^{+}$ and $H^{-}$, respectively. Without loss of generality, assume three triangles are bordered by the three $3$-cycles in $H^{+}$: $(v_{2},v_{3},v_{6}),(v_{4},v_{7},v_{8}),$ and $(v_{4},v_{5},v_{8})$. Therefore, two triangular faces are incident to edge $v_{4}v_{8}$ in $H^{+}$. However, a mixed $n$-gon is also incident to edge $v_{4}v_{8}$ in $H^{+}$. Thus, three faces are incident to edge $v_{4}v_{8}$ in $H^{+}$, which contradicts Fact 2. The second claim immediately follows from our argument. ∎
Proposition 6.7.
It is not possible for $G$ to $2$-cell embed on the torus where every mixed $n$-gon is a mixed quadrilateral.
Proof.
In this proof, we use the term face set in reference to the four possibilities given in Proposition 6.3. For sake of contradiction, suppose $G$ does $2$-cell embed on the torus where every mixed $n$-gon is a mixed quadrilateral. Then every $e\in E_{p}$ is incident to two mixed quadrilateral faces in this embedding. First, note the cycle sets $(a),(b),(d),(f),\text{ and }(g)$ include edge $v_{4}v_{8}$. Thus, embeddings corresponding to these cycle sets cannot have face sets 1, 2, and 3 from Proposition 6.3, since otherwise, we contradict Lemma 6.6. We now consider cycle sets (a)(g)(a)-(g) as separate cases, keeping in mind the face set restrictions already placed on some of these scenarios.
Case 1: Consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(a)$. As noted in the paragraph above, we only need to consider face set $4$ here, where $f_{3}=4$ and $f_{4}=14$. To avoid a contradiction with Lemma 6.6, a triangular face must border the cycle $(v_{2},v_{3},v_{6})$ in each $H$-fiber. There must also be a quadrilateral incident to edge $v_{3}v_{4}$ that borders a $4$-cycle in that $H$-fiber. However, this face must be adjacent to $v_{3}v_{6}$, and so, $v_{3}v_{6}$ lies on the border to three faces, which contradicts Fact 2. The exact same argument can be applied to cycle sets $(b),(f),$ and $(g)$.
Case 2: Next, consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(c)$. Assume a triangle is placed in the face bounded by the 3-cycle ($v_{2},v_{3},v_{6}$) in one of the $H$-fibers. There must also be a face incident to edge $v_{6}v_{7}$ that is contained in that $H$-fiber. However, this face must be adjacent to either $v_{6}v_{2}$ or $v_{6}v_{3}$, meaning $v_{6}v_{3}$ or $v_{6}v_{2}$ are incident to three faces, which is a contradiction. This implies we can have a maximum of two triangular faces per $H$ fiber, and so, face sets 1, 2, and 3 are not possible here. Thus, we can assume $f_{3}=4$ and $f_{4}=14$, with triangles bounded by $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in both $H$-fibers. Since all remaining faces must be quadrilaterals, edge $v_{4}v_{5}$ must be on the boundary of some face bound by a $4$-cycle in an $H$-fiber. However, this means edge $v_{4}v_{7}$ must be incident to three faces, which contradicts Fact 2.
Case 3: Next, consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(d)$. By Lemma 6.6, we see that $f_{3}=4$, $f_{4}=14$, and a triangle must be bound by the cycle $(v_{2},v_{3},v_{6})$ in each of the $H$-fibers. Then there must be a quadrilateral in each $H$-fiber that contains edge $v_{2}v_{1}$ on its border. However, this face must also contain edge $v_{2}v_{6}$ on its border, and so, $v_{2}v_{6}$ borders three faces, which contradicts Fact 2.
Case 4: Next, consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(e)$. Following the same argument given in Case 3, we see that only face set $4$ is possible where $f_{3}=4$, $f_{4}=14$, and triangles are bound by $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in both $H$-fibers. Thus, there must be a quadrilateral whose boundary is a $4$-cycle in an $H$-fiber that contains edge $v_{5}v_{1}$. However, this face must be adjacent to either $v_{5}v_{8}$ or $v_{5}v_{4}$, which implies that either $v_{5}v_{8}$ or $v_{5}v_{4}$ is adjacent to three faces, which contradicts Fact 2. ∎
The previous case only considers potential toroidal embeddings of $G$ where every mixed $n$-gon is a mixed quadrilateral. We now consider potential toroidal embeddings of $G$ where some mixed $n$-gon is not a mixed quadrilateral, and so, either a mixed pentagon or a mixed hexagon must appear in such embeddings. Up to exchanging the roles of $H^{\pm}$, any mixed pentagon $p$ must be bound by a $5$-cycle in $G$ that contains two adjacent edges in one $H$-fiber, two edges in $E_{p}$, and one edge in the other $H$-fiber. There are two types of mixed hexagons that can occur. Up to exchanging roles of $H^{\pm}$, one type of mixed hexagon, $h_{1}$, contains three adjacent edges in $H^{+}$, two edges in $E_{p}$, and one edge in $H^{-}$. In the other case, the mixed hexagon $h_{2}$ has two edges in $H^{+}$, two edges in $E_{p}$, and two edges in $H^{-}$.
Lemma 6.8.
It is not possible for $G$ to 2-cell embed on the torus with the following conditions:
  1. (1)
    Three distinct triangular faces are bound by the 3-cycles in one $H$ fiber.
  2. (2)
    Two distinct quadrilateral faces are bound by 4-cycles in the same $H$ fiber.
Proof.
For sake of contradiction, suppose we have an embedding of $G$ that meets conditions (1) and (2). Without loss of generality, assume conditions (1) and (2) are applied to $H^{+}$. Note that, there are five distinct $4$-cycles in $H^{+}$: $(v_{3},v_{4},v_{7},v_{6})$, $(v_{1},v_{2},v_{6},v_{7})$, $(v_{1},v_{7},v_{8},v_{5})$, $(v_{4},v_{5},v_{8},v_{7})$, and $(v_{1},v_{7},v_{4},v_{5})$. Furthermore, note that our embedding must admit eight mixed $n$-gons since there are eight edges in $E_{p}$, edges in $E_{p}$ are only bordered by mixed $n$-gons in such an embedding, and each mixed $n$-gon borders exactly two edges in $E_{p}$.
Consider an embedding with a quadrilateral bound by the 4-cycle $(v_{1},v_{7},v_{8},v_{5})$ in $H^{+}$. A mixed $n$-gon must be incident to the edge $e_{p}$ in $E_{p}$ whose vertices correspond with $v_{8}$ in both $H$-fibers. Then $e_{p}$ must also be incident to some edge, $e$, in $H^{+}$ incident to $v_{8}$. Since triangular faces are bound by the cycles $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in $H^{+}$, we see that $e$ is incident to three faces, which contradicts Fact 2. The same type of argument can be applied to an embedding with a quadrilateral bound by the $4$-cycle $(v_{4},v_{5},v_{8},v_{7})$.
Next, consider an embedding with a quadrilateral bound by the 4-cycle $(v_{1},v_{7},v_{4},v_{5})$ in $H^{+}$. Since triangular faces are placed incident to the cycles $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in $H^{+}$, edges $v_{4}v_{5}$, $v_{4}v_{7}$, and $v_{4}v_{8}$ are incident to exactly two faces just from these two triangles and this one quadrilateral. However, there must be a mixed $n$-gon incident to the edge in $E_{p}$ whose endpoints correspond to $v_{4}$ in each $H$-fiber. The previous statement forces this mixed $n$-gon to contain the edge $v_{4}v_{3}$ on its boundary. Further, this mixed $n$-gon cannot contain the edges $v_{4}v_{5}$, $v_{4}v_{7}$, and $v_{4}v_{8}$ in $H^{+}$ on its boundary, since otherwise, Fact 2 is violated. However, any such mixed $n$-gon with these properties will result in a local neighborhood around the $v_{4}$ vertex in $H^{+}$ of this embedding that is not homeomorphic to a disk, which contradicts Fact 1.
Thus, the two quadrilaterals that bound $4$-cycles in $H^{+}$ must be bound by $(v_{1},v_{2},v_{6},v_{7})$ and $(v_{3},v_{6},v_{7},v_{4})$, respectively. Consider the edge $e_{p}\in E_{p}$ whose endpoints correspond to $v_{6}$ in each $H$-fiber. This embedding must have a mixed $n$-gon that contains $e_{p}$ on its boundary along with an edge in $H^{+}$ incident to $v_{6}$. However, this implies that some edge incident $v_{6}$ is incident to three faces, which contradicts Fact 2. ∎
Proposition 6.9.
It is not possible for $G$ to $2$-cell embed on the torus where some mixed $n$-gon is a mixed pentagon or a mixed hexagon.
Proof.
For sake of contradiction, suppose that $G$ does $2$-cell embeds on the torus where some mixed $n$-gon is a mixed pentagon or a mixed hexagon. By Proposition 6.3, we see that only face sets (1), (2), or (3) could be used in such an embedding. As noted in the proof of Lemma 6.8, our embedding must admit eight mixed $n$-gons.
Case 1: First, suppose face set (1) was used for this embedding, and so, $f_{3}=6$, $f_{4}=11$, and $f_{6}=1$. Thus, there exists one mixed hexagon and seven mixed quadrilaterals in this embedding. Since $f_{3}=6$, all three $3$-cycles in both $H$-fibers are bound by triangles. Thus, the remaining four quadrilaterals in this embedding must each bound a $4$-cycle in one of the $H$-fibers. This implies that at least two quadrilateral faces are bound by $4$-cycles in one of the $H$-fibers. Thus, Lemma 6.8 implies that such an embedding is not possible.
Case 2: Now, suppose face set (2) was used for this embedding, and so, $f_{3}=6$, $f_{4}=10$, and $f_{5}=2$. First, all three $3$-cycles in each $H$-fiber must be bound by triangles. Such an embedding either has one mixed pentagon and seven mixed quadrilaterals or two mixed pentagons and six mixed quadrilaterals. Either way, there must exist at least three quadrilateral faces whose boundaries are $4$-cycles contained in the $H$-fibers. Thus, there must exist at least two quadrilateral faces bound by $4$-cycles in either $H^{+}$ or $H^{-}$. As a result, Lemma 6.8 shows that such an embedding is not possible.
Case 3: Finally, suppose face set (3) was used for this embedding, and so, $f_{3}=5$, $f_{4}=12$, and $f_{5}=1$. Thus, in one of the $H$-fibers, say $H^{+}$, all three $3$-cycles are bound by triangles. In $H^{-}$, exactly two of the three $3$-cycles are bound by triangles. We also must have eight mixed $n$-gons, seven of which are mixed quadrilaterals and then one mixed pentagon, $p$. This implies that there are five quadrilateral faces whose boundaries are $4$-cycles, where each $4$-cycle is contained in a single $H$-fiber. To make sure each of the $13$ edges in each $H$-fiber borders exactly two faces, we must have two quadrilaterals bound by $4$-cycles in $H^{+}$, three quadrilaterals bound by $4$-cycles in $H^{-}$, two edges of $p$ contained in $H^{+}$, and one edge of $p$ contained in $H^{-}$. The conditions on $H^{+}$ combined with Lemma 6.8 show that such an embedding is not possible, completing the proof. ∎
We can now justify that $G$ has genus two.
Theorem 6.10.
Let $G=H\boxempty P_{2}$. Then $\gamma(G)=2$.
Proof.
Suppose $G$ does $2$-cell embed on the torus. Then by Lemma 6.4, every $n$-gon is bound by an $n$-cycle in $G$. By Proposition 6.3, we can only have faces that are triangles, quadrilaterals, pentagons, or hexagons in any such embedding. Furthermore, any $e\in E_{p}$ does not lie on a $3$-cycle, and so, any such $e\in E_{p}$ must lie on the boundary of a quadrilateral, pentagon, or hexagon in this embedding, i.e., there must exist mixed $n$-gons in this embedding. However, Proposition 6.7 and Proposition 6.9, show that these scenarios are not possible. Thus, $\gamma(G)>1$.
To see that $\gamma(G)=2$, we explicitly construct an embedding of $G$ on $S_{2}$. Embed each $H$-fiber of $G=H\boxempty P_{2}$ on a $2$-sphere, following the embedding in Figure 5. For each embedding of $H$ on the $2$-sphere, cut out the interiors of three faces, specifically, the ones bound by cycles $(v_{1},v_{2},v_{3},v_{4},v_{5})$, $(v_{3},v_{4},v_{7},v_{6})$, and $(v_{4},v_{5},v_{8})$, respectively. Then connect corresponding open patches in each thrice punctured sphere by a tube. One can easily embed all the necessary $P_{2}$ fiber edges on these tubes, providing an embedding of $G$ on a surface of genus $2$, as needed. ∎
7. Conclusion
In light of Conjecture 5.4 and the work accomplished here, it remains to prove (or disprove) the following: if $G$ is $3$-connected and $G\boxempty P_{2}$ is toroidal, then $G$ must be OC. One might be able to tackle this question by using the list of $38$ minor-minimal NOC graphs given in [2]. More specifically, one would need an efficient method to check if $G\boxempty P_{2}$ is toroidal where $G$ is one these $38$ minor-minimal NOC graphs.

### v3

@@
-Toroidal Cartesian Products Where One Factor is $3$-Connected
-Abstract.
-In this paper, we provide a partial classification of Cartesian products of graphs that embed on the torus where one factor is $3$-connected.
-We show that if $G\boxempty H$ embeds on the torus and $G$ is $3$-connected, then $H$ must be a path on two vertices, $P_{2}$, and we provide significant evidence that $G$ must be outer-cylindrical.
-At the same time, we show that if $G$ is outer-cylindrical, then $G\boxempty P_{2}$ embeds on the torus.
-This all leads us to conjecture that if $G$ is $3$-connected then $G\boxempty H$ embeds on the torus if and only if $G$ is outer-cylindrical and $H=P_{2}$.
-We also determine the genera of $K_{4}\boxempty P_{3}$ and $H_{i}\boxempty P_{2}$ where $H_{i}$ is a planar, $3$-connected, non-outer-cylindrical graph derived from a wheel graph $W_{i}$ via two vertex splits and one edge addition.
-1. Introduction
+
+Abstract
+In this paper, we show that if $G$ is $3$-connected, then the Cartesian product of graphs $G\boxempty H$ embeds on the torus if and only if $G$ is outer-cylindrical and $H$ is a path on two vertices, $P_{2}$.
+As a by-product of our work, we also show that $K_{4}\boxempty P_{3}$ has genus two.
+1 Introduction
 Classifying embeddings of graphs on surfaces is a well-studied topic in topological graph theory.
-Here, we are interested in embeddings of a Cartesian product of two graphs $G$ and $H$, denoted $G\boxempty H$, which is defined to have vertex set $V(G)\times V(H)$ and a pair of vertices $(u,v),(u^{\prime},v^{\prime})\in V(G)\times V(H)$ determine an edge in $G\boxempty H$ if and only $u=u^{\prime}$ and $vv)\in E(H)$ or $uu^{\prime}\in E(G)$ and $v=v^{\prime}$.
-We refer the reader to [7] for further background on graph Cartesian products.
-The fiber structures of graph Cartesian products can sometimes be leveraged to efficiently build embeddings and determine genera of such graphs; see [10], [11], [13], [14], [15], [17], [20], [21] for some examples.
+Here, we are interested in embeddings of a Cartesian product of two graphs $G$ and $H$, denoted $G\boxempty H$, which is defined to have vertex set $V(G)\times V(H)$ and a pair of vertices $(u,v),(u^{\prime},v^{\prime})\in V(G)\times V(H)$ determine an edge in $G\boxempty H$ if and only if $u=u^{\prime}$ and $vv^{\prime}\in E(H)$ or $uu^{\prime}\in E(G)$ and $v=v^{\prime}$.
+We refer to $G$ and $H$ as the factors of $G\boxempty H$.
+Every $G\times\{h\}$ with $h\in V(H)$ is called a $G$-fiber of $G\boxempty H$, and $H$-fibers are defined analogously.
+We refer the reader to ImKlRa2008 for further background on graph Cartesian products.
+The fiber structures of graph Cartesian products can sometimes be leveraged to efficiently build embeddings and determine genera of such graphs; see MiSa2022 , MoPiWh1990 , Pi1980 , Pi1982 , Pi1989 , Sun2023 , Wh1970 , Wh1971 for some examples.
 In particular, it would be interesting to know if one can determine the set of all Cartesian products of graphs that embed on a fixed surface.
-Such a classification was established for planar graphs, that is, graphs that embed in the plane, in the work of Behzad and Mahmodian [4].
-More recently, the second author along with Abell and McDermott have provided a complete classification of Cartesian products of graphs that embed in the projective plane in [1].
+Such a classification was established for planar graphs, that is, graphs that embed in the plane, in the work of Behzad and Mahmoodian BeMa1969 .
+More recently, the second author along with Abell and McDermott have provided a complete classification of Cartesian products of graphs that embed in the projective plane in AbMcMi2025 .
 In this paper, we are interested in classifying which Cartesian product of graphs are toroidal, i.e., embed on the torus.
 While the classifications of planar and projective planar Cartesian products were reasonably straightforward, we expect the classification of toroidal Cartesian products to be far more complicated.
-To start, the work of Kuratowski [8] and Wagner [19] showed there are only two minor-minimal obstructions to planarity, namely $K_{5}$ and $K_{3,3}$, and Archdeacon in [3] showed that there is a complete list of 35 graphs that are minor-minimal obstructions to embedding on the projective plane.
-However, the list of minor-minimal obstructions for embedding on the torus isn’t completely known, and this list contains at least 17 thousand forbidden minors; see [12].
+To start, the work of Kuratowski Ku1930 and Wagner Wa1937 showed there are only two minor-minimal obstructions to planarity, namely $K_{5}$ and $K_{3,3}$, and Archdeacon in Ar1980 showed that there is a complete list of 35 graphs that are minor-minimal obstructions to embedding on the projective plane.
+However, the list of minor-minimal obstructions for embedding on the torus is not completely known, and this list contains at least 17 thousand forbidden minors; see MyWo2018 .
 In addition, when restricting to Cartesian products of graphs, we see far more flexibility in toroidal embeddings than planar or projective planar embeddings with regards to (vertex) connectivity.
 If $G\boxempty H$ is planar or projective planar, then both $G$ and $H$ are each at most $2$-connected.
-This follows from the fact that $K_{4}\boxempty P_{2}$ is both non-planar and non-projective planar, and any $3$-connected graph contains a $K_{4}$ minor; see Lemma 4.2.
-However, in Section 5, we provide infinite classes of $3$-connected and $4$-connected graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal.
+This follows from the fact that $K_{4}\boxempty P_{2}$ is both non-planar and non-projective planar, and any $3$-connected graph contains a $K_{4}$ minor; see Lemma 3.
+However, Example 1 in Section 5 shows that there exist $G\boxempty P_{2}$ that are toroidal where $G$ is $4$-connected.
 Thus, even a partial classification of toroidal Cartesian products could be a far more complicated task than the planar and projective planar cases.
 Here, we provide some major first steps towards this classification problem, with the additional assumption that one factor is $3$-connected.
-A simple combinatorial argument shows that if $G\boxempty H$ is toroidal, then $G$ and $H$ can each be at most $4$-connected in Proposition 4.1.
-Then in Theorem 4.3 we show that if $G$ is $3$-connected, then any toroidal $G\boxempty H$ must have $H=P_{2}$.
+In what follows, a graph $G$ is outer-cylindrical if there exists a planar embedding of $G$ such that the vertices of $G$ all lie on the boundary of at most two faces in this embedding.
+Theorem 1.1.
+Let $H$ and $G$ be simple, connected, nontrivial graphs, and suppose $G$ is $3$-connected. Then $G\boxempty H$ is toroidal if and only if $G$ is an outer-cylindrical graph and $H=P_{2}$.
+The proof of Theorem 1.1 is broken up into a few pieces.
+In Proposition 5, we show that if $G$ is $3$-connected, then any toroidal $G\boxempty H$ must have $H=P_{2}$.
 This proof follows from the fact that $K_{4}\boxempty P_{3}$ is not toroidal, which is proved in Theorem 3.1.
-In Section 5, we provide two explicit constructions of infinite classes of toroidal $G_{i}\boxempty P_{2}$ and observe that any such $G_{i}$ in these constructions is planar and outer-cylindrical; see the beginning of Section 5 for the definition of outer-cylindrical.
-This all leads us to Conjecture 5.4, which states that if $G$ is $3$-connected, then $G\boxempty H$ is toroidal if and only if $H=P_{2}$ and and $G$ outer-cylindrical.
-Our work proves the backwards direction of this conjecture and provides strong evidence in favor of the forward direction.
-Theorem 4.3 already shows that $H=P_{2}$ holds under the hypotheses of the forward direction.
-Section 6 provides further evidence that the forward direction of this conjecture holds by showing the existence of an infinite set of planar, $3$-connected, non-outer-cylindrical graphs $\{H_{i}\}$ such that each $H_{i}\boxempty P_{2}$ does not embed on the torus and $H_{i}$ is “not far” in some sense from a wheel graph, $W_{i}$, which is an outer-cylindrical $3$-connected graph.
-There are two significant by-products of our work on this classification problem.
-In Section 3 we prove that the genus of $K_{4}\boxempty P_{3}$ is two, highlighted in Theorem 3.4.
-This work requires a careful combinatorial analysis of all potential embeddings of this graph on a torus, which employs techniques similar to [5].
+Proposition 6 provides an explicit construction for building a toroidal embedding of $G\boxempty P_{2}$, where $G$ is outer-cylindrical.
+Then Proposition 7 shows that if $G\boxempty P_{2}$ is toroidal, then $G$ must be an outer-cylindrical graph.
+This proof only requires some basic combinatorial topology of graph embeddings.
+A significant by-product of our work occurs in Section 3 where we prove that the genus of $K_{4}\boxempty P_{3}$ is two.
+This work requires a careful combinatorial analysis of all potential embeddings of this graph on a torus, which employs techniques similar to BrSq1988 .
 In particular, even though the girth of $K_{4}\boxempty P_{3}$ is three, one cannot hope for a triangular embedding since every edge in a $P_{3}$ fiber does not lie on a $3$-cycle.
-Thus, determining the genera of such a graph is not a straightforward task.
-In addition, in Section 6, we prove Theorem 6.10, which shows that $H\boxempty P_{2}$ has genus two, where $H$ is a specified planar, $3$-connected, non-outer-cylindrical graph.
-Parts of this proof follow the structure of the proof of Theorem 3.4, but the analysis is much more complex.
-As far as the authors can tell, both Theorem 3.4 and Theorem 6.10 provide novel contributions to the topological graph theory literature.
-We would like to acknowledge the Furman University Mathematics Department for financially supporting this work via the Summer Mathematics Undergraduate Research Fellowships.
-2. Background
+Thus, determining the genus of such a graph is not a straightforward task.
+2 Background
 In this paper, we assume all graphs are simple, connected, and non-trivial (in the sense of being nonempty and not a single vertex), unless otherwise noted.
 We use $S_{g}$ to denote an orientable, closed, connected surface of genus $g\in\mathbb{N}\cup\{0\}$.
-Given an embedding of a graph $G$ on a surface $S_{g}$ via a function $f:|G|\rightarrow S_{g}$ where $|G|$ represents a geometric realization of $G$, we sometimes use $f(G)$ to denote this embedding or just $G$ when the embedding is clear from context.
+Given an embedding of a graph $G$ on a surface $S_{g}$ via a continuous injective function $f:|G|\rightarrow S_{g}$ where $|G|$ represents a geometric realization of $G$, we sometimes use $f(G)$ to denote this embedding or just $G$ when the embedding is clear from context.
 If each component of $S_{g}\setminus f(G)$ is homeomorphic to an open disk, then we say this embedding is a $2$-cell embedding.
@@
 An embedding of $G$ that realizes the minimal genus is called a minimal embedding.
-We will mainly be interested in $2$-cell embeddings of graphs, and since the work of Youngs [22] shows that each minimal embedding of $G$ is a $2$-cell embedding, we usually don’t need to make a distinction when determining the genera of a graph.
-We refer the reader to [6] for further background on graph embeddings on surfaces.
+We will mainly be interested in $2$-cell embeddings of graphs, and since the work of Youngs Yo1963 shows that each minimal embedding of $G$ is a $2$-cell embedding, we usually do not need to make a distinction when determining the genera of a graph.
+We refer the reader to GrTu2001 for further background on graph embeddings on surfaces.
 A $2$-cell embedding $f:|G|\rightarrow S_{g}$ provides a $2$-cell decomposition of $S_{g}$ with $|V(G)|$ vertices, $|E(G)|$ edges, and $|F|$ faces, where each face is a component of $S_{g}\setminus f(G)$.
@@
 Recall that the girth of a graph $G$ (that is not a tree) is the length of its shortest cycle.
-Proposition 2.1.
+Proposition 1.
 Let $G$ be a graph (that is not a tree) with girth $\alpha$. Then
@@
 We note the following basic but important facts about $2$-cell embeddings of $G$:
-Fact 1: Each vertex of $G$ has an open $2$-cell neighborhood in $S$.
-Fact 2: Each edge of $G$ borders precisely two (not necessarily unique) faces.
+Fact 1: Each vertex of $G$ has an open $2$-cell neighborhood in $S_{g}$.
+Fact 2: Each edge of $G$ is either incident to two distinct faces or twice incident to one face.
 Fact 3: If $G=H\boxempty P_{n}$ for some graph $H$, then each edge from a $P_{n}$ fiber borders two $k$-gon faces, where $k\geq 4$.
 Fact 4: $2|E(G)|=\displaystyle\sum_{i=3}^{\infty}if_{i}$.
-Facts (1) and (2) are general facts about $2$-cell embeddings.
-Fact (3) follows from (2) and since any edge in a $P_{n}$ fiber is not contained in a $3$-cycle.
-Fact (4) follows from (2).
-This collection of facts will play an important role in ruling out potential toroidal embeddings of graphs in Section 3 and Section 6.
+Facts 1 and 2 are general facts about $2$-cell embeddings.
+Fact 3 follows from Fact 2 and since any edge in a $P_{n}$ fiber is not contained in a $3$-cycle.
+Fact 4 follows from Fact 2.
+This collection of facts will play an important role in ruling out potential toroidal embeddings of graphs in Section 3.
 A graph $G^{\prime}$ is a minor of $G$ if $G^{\prime}$ can be obtained by a series of vertex deletions, edge deletions, or edge contractions on $G$.
-Another way to find a lower bound for the genus of a graph $G$ is by finding the genus (or lower bound on the genus) for a subgraph or minor $G$.
+Another way to find a lower bound for the genus of a graph $G$ is by finding the genus (or lower bound on the genus) for a subgraph or minor of $G$.
 Since we will be interested in Cartesian products of graphs in this paper, we first confirm that the product of minors from each factor results in a minor of the product.
-Proposition 2.2.
+Proposition 2.
 If $G^{\prime}$ and $H^{\prime}$ are minors of $G$ and $H$ then $G^{\prime}\boxempty H^{\prime}$ is a minor of $G\boxempty H$.
@@
 Let $G$ and $H$ be graphs and suppose $G^{\prime}$ be a minor of $G$.
-Recall that the vertex sets and edge sets of $G\boxempty H$ are defined as $V(G\boxempty H)=V(G)\times V(H)$ and $E(G\boxempty H)=(V(G)\times E(H))\cup(E(G)\times V(H))$.
+Recall that the vertex sets and edge sets of $G\boxempty H$ are defined as $V(G\boxempty H)=V(G)\times V(H)$ and $E(G\boxempty H)=\{\{(u,v),(u,v^{\prime})\}:u\in V(G),vv^{\prime}\in E(H)\}\cup\{\{(u,v),(u^{\prime},v)\}:uu^{\prime}\in E(G),v\in V(H)\}$.
 It suffices to show $G^{\prime}\boxempty H$ is a minor of $G\boxempty H$ under an edge deletion, vertex deletion, or edge contraction, which we now handle as three separate cases.
-Case 1: Let $G^{\prime}=G-e$ for some $e\in E(G)$. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by deleting edge $(e,v)\in E(G)\times V(H)$, for each $v\in V(H)$.
+Case 1: Let $G^{\prime}=G-e$ for some $e=uu^{\prime}\in E(G)$.
+Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by deleting edge $\{(u,v),(u^{\prime},v)\}\in E(G\boxempty H)$, for each $v\in V(H)$.
 Case 2: Let $G^{\prime}=G-v_{0}$ for some $v_{0}\in V(G)$.
@@
 Then we can obtain $G^{\prime}\boxempty H$ from $G\boxempty H$ by deleting vertex $(v_{0},v)\in V(G)\times V(H)$, for each $v\in V(H)$ (and deleting all its incident edges).
-Case 3: Given edge $e\in E(G)$ with endpoints $a$ and $b$, we edge contract $e$ in $G$ to create $G^{\prime}$ which now has a new vertex $u\in V(G^{\prime})$ to which $a$ and $b$ have been identified.
-Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by contracting edge $(e,v)\in E(G)\times V(H)$, for each $v\in V(H)$.
+Case 3: Given edge $e=uu^{\prime}\in E(G)$, we contract the edge $e$ in $G$ to create $G^{\prime}$ which now has a new vertex $a\in V(G^{\prime})$ to which $u$ and $u^{\prime}$ have been identified.
+Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by contracting edge $\{(u,v),(u^{\prime},v)\}\in E(G\boxempty H)$, for each $v\in V(H)$.
 ∎
-Corollary 2.3.
+Corollary 1.
 Suppose $G^{\prime}$ and $H^{\prime}$ are minors (or subgraphs) of graphs $G$ and $H$, respectively. Then $\gamma(G\boxempty H)\geq\gamma(G^{\prime}\boxempty H^{\prime})$.
-3. The genus of $K_{4}\boxempty P_{3}$
+3 The genus of $K_{4}\boxempty P_{3}$
 In this section, we will show that $\gamma(K_{4}\boxempty P_{3})=2$.
-Since every $3$-connected graph contains a $K_{4}$ minor (see Lemma 4.2), this will play a pivotal role in classifying toroidal Cartesian products where one of the factors is $3$-connected.
+Since every $3$-connected graph contains a $K_{4}$ minor (see Lemma 3), this will play a pivotal role in classifying toroidal Cartesian products where one of the factors is $3$-connected.
 For the remainder of this section, set $G=K_{4}\boxempty P_{3}$.
@@
 In particular, no edge contained in a $P_{3}$ fiber is contained in a $3$-cycle.
-Lemma 3.1.
+Lemma 1.
 For any $2$-cell embedding of $G$, at most two $3$-cycles in each $K_{4}$ fiber are triangles, and so, $f_{3}\leq 6$.
@@
 Our main goal is to now prove that $\gamma(G)\neq 1$ by showing that an embedding of $G$ with $|F|=14$ is not possible.
-Proposition 3.2.
+Proposition 3.
 Suppose $G$ admits a $2$-cell embedding on the torus. Then this embedding has $|F|=14$ with the following possibilities:
-  1. (1)
+  1. 1.
     $f_{3}=4$ and $f_{4}=10$,
-  2. (2)
+  2. 2.
     $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$,
-  3. (3)
+  3. 3.
     $f_{3}=6$, $f_{4}=6$, and $f_{5}=2$, or
-  4. (4)
+  4. 4.
     $f_{3}=6$, $f_{4}=7$, and $f_{6}=1$.
@@
 We will frequently make use of this type of combinatorial restriction in this proof, though we will suppress most of the computations moving forward.
-Combined with Lemma 3.1, we have $4\leq f_{3}\leq 6$.
+Combined with Lemma 1, we have $4\leq f_{3}\leq 6$.
 First, suppose $f_{3}=6$.
@@
 Finally, suppose $f_{3}=4$. Then $\sum_{i=4}^{\infty}f_{i}=10$ and $\sum_{i=4}^{\infty}if_{i}=40$, which can only occur if all ten faces are quadrilaterals. ∎
-In light of Proposition 3.2, it will help to more carefully describe how faces can behave in a $2$-cell embedding of $G$ on the torus.
+In light of Proposition 3, it will help to more carefully describe how faces can behave in a $2$-cell embedding of $G$ on the torus.
 We first introduce some necessary notation.
@@
 We use $E_{1}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{1}$ and $K^{2}$ and $E_{2}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{2}$ and $K^{3}$.
-Given a $2$-cell embedding of $G$, we say that a quadrilateral is a mixed quadrilateral if it’s boundary is a $4$-cycle of $G$ that consists of two edges from $E_{1}$ ($E_{2}$), one edge from a $K^{1}$ ($K^{3}$), and the corresponding edge from $K^{2}$.
+Given a $2$-cell embedding of $G$, we say that a quadrilateral is a mixed quadrilateral if its boundary is a $4$-cycle of $G$ that consists of two edges from $E_{1}$ ($E_{2}$), one edge from a $K^{1}$ ($K^{3}$), and the corresponding edge from $K^{2}$.
 Note, that any quadrilateral in an embedding of $G$ that contains a $P_{3}$ fiber edge must be a mixed quadrilateral.
@@
 The 6-cycle boundary of $h_{3}$ contains one edge in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K_{4}$.
-The $6$-cycle boundary of $h_{4}$ contains one edge in an outer $K_{4}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K^{2}$.
+The $6$-cycle boundary of $h_{4}$ contains one edge in an outer $K_{4}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in $K^{2}$.
 For the above description, we only considered faces in a $2$-cell embedding of $G$ on the torus whose boundary was an appropriate cycle.
 We now show those are the only types of faces that can occur in such an embedding.
-Lemma 3.3.
-If $G$ has a $2$-cell embedding on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle of $G$.
+Lemma 2.
+If $G$ admits a $2$-cell embedding on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle of $G$.
 Proof.
@@
 Then the boundary of $F$ is a closed walk in $G$ determined by a “facial walk” of this embedding.
-From Proposition 3.2, we only need to consider when $F$ is a triangle, quadrilateral, pentagon, or hexagon.
-We will show that if the facial walk for $F$ isn’t an appropriate cycle, then we always reach a contradiction.
+From Proposition 3, we only need to consider when $F$ is a triangle, quadrilateral, pentagon, or hexagon.
+We will show that if the facial walk for $F$ is not an appropriate cycle, then we always reach a contradiction.
 First, suppose our embedding admits an $n$-gon face, $F$, that is not bound by an $n$-cycle for some $n=3$, 44, or $5$.
-The boundary of $F$ is must be a closed walk in $G$, and since the girth of $G$ is three, this closed walk must have a subwalk of the form eveeve for some $e\in E(G)$ and $v\in V(G)$.
+The boundary of $F$ must be a closed walk in $G$, and since the girth of $G$ is three, this closed walk must have a subwalk of the form eveeve for some $e\in E(G)$ and $v\in V(G)$.
 However, this implies that $G$ admits a degree one vertex, which is a contradiction.
-Now, suppose $F$ is a hexagon and the facial walk of $F$ determines a closed walk of length six in $G$ that isn’t a $6$-cycle.
-If the facial walk for $F$ doesn’t contain any cycle subgraphs, then we once again get a degree one vertex in $G$, which is a contradiction.
+Now, suppose $F$ is a hexagon and the facial walk of $F$ determines a closed walk of length six in $G$ that is not a $6$-cycle.
+If the facial walk for $F$ does not contain any cycle subgraphs, then we once again get a degree one vertex in $G$, which is a contradiction.
 However, in any $K_{4}$ fiber of $G$, there exist pairs of $3$-cycles that intersect along a single edge.
@@
 However, such an $F$ is not possible.
-If $f_{6}\geq 1$, then Proposition 3.2 shows that $f_{3}=6$ and $f_{4}=7$.
+If $f_{6}\geq 1$, then Proposition 3 shows that $f_{3}=6$ and $f_{4}=7$.
 Note that, any triangle in this embedding can only bound edges in $K_{4}$ fibers, and we just assumed our one hexagon only bounds edges in $K_{4}$ fibers.
@@
 Proof.
-We will first show that $G=K_{4}\boxempty P_{3}$ does not embed on the torus by analyzing the four possibilities given in Proposition 3.2. We then show that $G$ does admit an embedding on $S_{2}$.
-For sake of contradiction, suppose $G$ 22-cell embeds on the torus with $|F|=14$.
-In each case, we show that our assumpotions violate Fact (2), giving a necessary contradiction.
+We will first show that $G=K_{4}\boxempty P_{3}$ does not embed on the torus by analyzing the four possibilities given in Proposition 3. We then show that $G$ does admit an embedding on $S_{2}$.
+Suppose $G$ admits a $2$-cell embedding on the torus with $|F|=14$.
+In each case, we show that our assumptions violate Fact (2), giving a necessary contradiction.
 The following is an important fact we will use throughout the proof: any $e\in E_{1}\cup E_{2}$ does not lie on a $3$-cycle and if such an edge $e$ is bound by a quadrilateral in this embedding, then it must be a mixed quadrilateral.
@@
 At this point, all faces incident to edges in $K^{2}$ have been determined.
-Thus, by Lemma 3.1 and our assumptions on $f_{3}$ and $f_{4}$, there are two triangles incident to only edges in $K^{1}$, two triangles incident to only edges in $K^{3}$, and one quadrilateral that must be incident to either only edges in $K^{1}$ or $K^{3}$.
+Thus, by Lemma 1 and our assumptions on $f_{3}$ and $f_{4}$, there are two triangles incident to only edges in $K^{1}$, two triangles incident to only edges in $K^{3}$, and one quadrilateral that must be incident to either only edges in $K^{1}$ or $K^{3}$.
 Note that $K^{1}$ and $K^{3}$ each have four edges incident to mixed quadrilaterals.
-Thus, either some edges in $K^{1}$ don’t border two faces and some edges in $K^{2}$ border at least three faces, or vice versa, giving a contradiction.
-Suppose $|F|=14$ and $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$.
+Thus, either some edges in $K^{1}$ do not border two faces and some edges in $K^{2}$ border at least three faces, or vice versa, giving a contradiction.
+Suppose $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$.
 The two types of pentagons ($p_{1}$ and $p_{2}$ described earlier in this section) each bound two edges coming from either $E_{1}$ or $E_{2}$.
@@
 Since each mixed quadrilateral is incident to an edge in $K^{2}$ and a pentagon is incident to either one or two edges in $K^{2}$, either four or three more (not necessarily distinct) edges in $K^{2}$ must be incident to the other faces in this embedding.
-By Lemma 3.1 and $f_{3}=5$, either one or two triangles are incident to only edges in $K^{2}$.
-Having two triangles contained in $K^{2}$ results in some edge in $K^{2}$ having at least three faces incident to it, which is a contradiction.
-Having only one triangle contained in $K^{2}$ results in some edge in $K^{2}$ having only one face incident to it, which is a contradiction.
-The only other possibility is that one triangle and the one (non-mixed) quadrilateral are incident to edges in $K^{2}$, but this would again result in some edge of $K^{2}$ having three faces incident to it.
+By Lemma 1 and $f_{3}=5$, either one or two triangles are incident to only edges in $K^{2}$.
+Having two triangles bordered by edges in $K^{2}$ results in some edge in $K^{2}$ having at least three faces incident to it, which is a contradiction.
+So, suppose there is only one triangle bordered by edges in $K^{2}$, and so, each outer $K_{4}$ is bordered by two triangles.
+In this case, a $p_{2}$ pentagon must occur to make sure each edge of $K^{2}$ is bordered by exactly two faces.
+This also means the one remaining quadrilateral must be incident to only edges in a single outer $K_{4}$, say $K^{1}$.
+Then some edge in $K^{1}$ has three faces incident to it.
 Thus, an embedding of $G$ with $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$ is not possible.
 Suppose $f_{3}=6$, $f_{4}=6$, and $f_{5}=2$.
-By Lemma 3.1, each $K_{4}$ fiber contains exactly two triangles.
+By Lemma 1, each $K_{4}$ fiber contains exactly two triangles.
 Note that, each pentagon borders a pair of edges in either $E_{1}$ or $E_{2}$.
@@
 Suppose $f_{3}=6$, $f_{4}=7$, and $f_{6}=1$.
-By Lemma 3.1, each $K_{4}$ fiber contains exactly two triangles.
-As discussed before the proof, there are three types of hexagons that can exist in an embedding of $G$.
+By Lemma 1, each $K_{4}$ fiber contains exactly two triangles.
+As discussed before the proof, there are four types of hexagons that can exist in an embedding of $G$.
 We now break this case into sub-cases, depending on what type of hexagon is used.
@@
 The four edges of $E_{1}$ can be embedded on these two tubes.
-Now, delete the triangular faces $F_{3}$ and $F_{4}$ coming from $K^{2}$’s embedding on a $2$-sphere and also delete $F_{3}$ and $F_{4}$ from $K^{3}$’s embedding on a $2$-sphere.
+Now, delete the triangular faces $F_{3}$ and $F_{4}$ coming from the $K^{2}$ embedding on a $2$-sphere and also delete $F_{3}$ and $F_{4}$ from the $K^{3}$ embedding on a $2$-sphere.
 Insert two tubes in a similar fashion to embed the edges of $E_{2}$, completing the embedding of $G$.
@@
 ∎
-4. Restricting toroidal $G\boxempty H$ via vertex connectivity.
+4 Restricting toroidal $G\boxempty H$ via vertex connectivity.
 The vertex-connectivity of a graph $G$, denoted $\kappa(G)$, is the minimum cardinality of a vertex-cut if $G$ is not complete or $\kappa(G)=n-1$ if $G=K_{n}$ for $n\in\mathbb{N}$.
@@
 We first note that $G$ and $H$ can be at most $4$-connected, if $G\boxempty H$ is toroidal.
-Proposition 4.1.
+Proposition 4.
 If $\gamma(G\boxempty H)\leq 1$, then $\kappa(G),\kappa(H)\leq 4$.
@@
 Since $G$ is simple, the girth of $G\boxempty P_{2}$ is at least three.
-By Proposition 2.1, we have that
+By Proposition 1, we have that
 $\gamma(G\boxempty P_{2})\geq 1+\frac{2|E(G)|+|V(G)|}{6}-\frac{2|V(G)|}{2}=1+\frac{1}{3}|E(G)|-\frac{5}{6}|V(G)|.$
-Since $G$ is $5$-connected, for all $v\in V(G)$, we have $deg(v)\geq 5$.
-Since $G$ is simple, every edge is incident to two distinct vertices.
-Thus, $\frac{5}{2}|V(G)|\leq|E(G)|$, and so, we can conclude that
-$\gamma(G\boxempty P_{2})\geq 1+\frac{1}{3}(\frac{5}{2}|V(G)|)-\frac{5}{6}|V(G)|=1.$
+Since $G$ is $5$-connected, for all $v\in V(G)$, we have $deg(v)\geq 5$. By the Handshaking Lemma, $\sum deg(v)=2|E(G)|$. Thus, $\frac{5}{2}|V(G)|\leq|E(G)|$, and so, we can conclude that
+$\gamma(G\boxempty P_{2})\geq 1+\frac{1}{3}\left(\frac{5}{2}|V(G)|\right)-\frac{5}{6}|V(G)|=1.$
 Equality holds if and only if $G$ is $5$-regular and $G\boxempty P_{2}$ admits a triangular emdedding.
 However, any edge in a $P_{2}$ fiber is not contained in any $3$-cycle of $G\boxempty P_{2}$, and so, a triangular embedding of $G\boxempty P_{2}$ does not exist.
-Therefore, $\gamma(G\boxempty H)\geq\gamma(G\boxempty P_{2})>1.$
+Corollary 1 implies that $\gamma(G\boxempty H)\geq\gamma(G\boxempty P_{2})>1.$
 We now work to show that severe restrictions can be placed on $H$ when $\kappa(G)\in\{3,4\}$. We first need a basic fact about $3$-connected graphs.
-Lemma 4.2.
+Lemma 3.
 Let $G$ be a $3$-connected graph. Then $G$ contains a $K_{4}$ minor.
 Proof.
-The work of Tutte [18] shows that any simple, $3$-connected graph contains a wheel graph $W_{n}$ as a minor, for some $n\geq 3$; see Figure 2 for a diagram of $W_{5}$.
-At the same time, $W_{m}$ is a minor of $W_{n}$, whenever $m\leq n$.
-It is easy to see that one can first delete an appropriate number of the spoke edges of $W_{n}$ that connect to the central vertex, and then, contract some of the edges along the $n$-cycle of $W_{n}$ to obtain a $W_{m}$ minor.
-Thus, $W_{3}=K_{4}$ is a minor of any simple, $3$-connected graph, as needed.
+The work of Tutte Tu1961 shows that any simple $3$-connected graph contains a wheel graph $W_{n}$ as a minor, for some $n\geq 3$.
+Whenever $m\leq n$, it is easy to see that $W_{m}$ is a minor of $W_{n}$: one can first delete an appropriate number of the spoke edges of $W_{n}$ that connect to the central vertex, and then, contract some of the edges along the $n$-cycle of $W_{n}$ to obtain a $W_{m}$ minor.
+Thus, $W_{3}=K_{4}$ is a minor of any simple $3$-connected graph, as needed.
 ∎
 This result has immediate consequences for classifying toroidal Cartesian products where one factor is $3$-connected.
-Theorem 4.3.
+Proposition 5.
 Let $G$ be a $3$-connected graph. If $\gamma(G\boxempty H)\leq 1$, then $H=P_{2}$.
@@
 Then $H$ contains a $P_{3}$ subgraph.
-By Lemma 4.2, $G$ has a $K_{4}$ minor.
-Then by Theorem 3.4 and Corollary 2.3, we have that $\gamma(G\boxempty H)\geq\gamma(K_{4}\boxempty P_{3})=2$.
+By Lemma 3, we know that $G$ has a $K_{4}$ minor.
+Then by Theorem 3.1 and Corollary 1, we have that $\gamma(G\boxempty H)\geq\gamma(K_{4}\boxempty P_{3})=2$.
 ∎
-Combined, Theorem 4.3 and Proposition 4.1 give rise to the following questions:
-Question 4.4.
-Does there exist an infinite set of $3$-connected (44-connected) graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal? If so, can we classify such $G_{i}$?
-5. Constructing toroidal $G\boxempty P_{2}$, where $\kappa(G)\in\{3,4\}$.
-In this section, we address the first part of Question 4.4 and provide a conjecture for the second part of Question 4.4.
-This first requires us to introduce a new class of graphs.
-A graph $G$ is outer-cylindrical if it embeds on the $2$-sphere such that there are two distinct faces whose boundaries collectively contain all the vertices of $G$.
-Since the class of outer-cylindrical graphs is closed under taking minors, the Graph Minor Theorem of Robertson and Seymour [16] implies that there is a finite list of minor-minimal obstructions for this class of graphs, which were then determined by Archdeacon–Bonnington–Dean–Hartsfield–Scott [2].
-Following their notation, we will use the abbreviation OC to stand for outer-cylindrical and NOC to stand for non-outer-cylindrical.
-The following proposition motivates our interest in OC graphs.
-Proposition 5.1.
-Let $G$ be an OC-graph. Then $\gamma(G\boxempty P_{2})\leq 1$.
+Combined, Proposition 5 and Proposition 4 give rise to the following questions: Does there exist an infinite set of 33-connected (44-connected) graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal?
+If so, can we classify such $G_{i}$?
+5 Classifying toroidal $G\boxempty P_{2}$, where $\kappa(G)\in\{3,4\}$.
+In this section, we fully address both of questions from above.
+Recall that a graph $G$ is outer-cylindrical if it embeds in the plane or the $2$-sphere such that there are two distinct faces whose boundaries collectively contain all the vertices of $G$.
+Since the class of outer-cylindrical graphs is closed under taking minors, the Graph Minor Theorem of Robertson and Seymour RoSe2004 implies that there is a finite list of minor-minimal obstructions for this class of graphs, which were then determined by Archdeacon–Bonnington–Dean–Hartsfield–Scott ArBo2001 .
+The following proposition motivates our interest in outer-cylindrical graphs.
+Proposition 6.
+If $G$ is an outer-cylindrical graph, then $\gamma(G\boxempty P_{2})\leq 1$.
 Proof.
-We first construct disjoint OC-embeddings of the two $G$ fibers (call them $G_{1}$ and $G_{2}$) of $G\boxempty P_{2}$ in the $2$-sphere $\mathbb{R}\cup\{\infty\}$.
-Take an OC-embedding of $G$ as $G_{1}$, and reflect $G_{1}$ over a line that does not intersect $G_{1}$ to build the OC-embedding of $G_{2}$.
+We first construct disjoint outer-cylindrical embeddings of the two $G$ fibers (call them $G_{1}$ and $G_{2}$) of $G\boxempty P_{2}$ in the $2$-sphere.
+Take an outer-cylindrical embedding of $G$ as $G_{1}$, and reflect $G_{1}$ over a line that does not intersect $G_{1}$ to build the outer-cylindrical embedding of $G_{2}$.
 All the vertices of each $G_{i}$ for $i=1,2$ are contained in two faces of this embedding, and one of these faces, which we call the outer face, is common between $G_{1}$ and $G_{2}$.
 For each vertex $v$ in $G_{1}$ on this outer face, we can embed on edge on the $2$-sphere connecting $v$ to the corresponding vertex in $G_{2}$.
-We can then cut out the interiors of the “inner” faces for $G_{1}$ and $G_{2}$ and add a tube between them carrying the remaining edges of $G\boxempty P_{2}$ to build an embedding of $G\boxempty P_{2}$ on the torus.∎
+We can then cut out the interiors of the “inner” faces for $G_{1}$ and $G_{2}$ and add a tube between them carrying the remaining edges of $G\boxempty P_{2}$ to build an embedding of $G\boxempty P_{2}$ on the torus.
+See Figure 1 for a schematic of this embedding.
+∎
 […]
-Figure 1.
-The inner and outer faces for the $G_{1}$ fiber are depicted in green and the inner and outer faces for the $G_{2}$ fiber are depicted in black.
+Figure 1: The inner and outer faces for the $G_{1}$ fiber are depicted in green and the inner and outer faces for the $G_{2}$ fiber are depicted in black.
 The edges contained in the $P_{2}$ fibers that connect to vertices on the two outer faces are depicted in blue.
 The edges contained in the $P_{2}$ fibers that connect to vertices on the inner faces are not depicted.
-It is easy to see that there exist infinite classes of $3$-connected graphs that are OC.
-For this, we recall the definition of a wheel of order n for $n\geq 3$, denoted $W_{n}$, as the simple connected graph with $n+1$ vertices and $2n$ edges with the following properties:
-  1. (1)
-    $V(W_{n})=\{c,b_{0},\ldots,b_{n-1}\}$,
-  2. (2)
-    $c$ is adjacent to each $b_{i}$ for $i=0,\ldots n-1$,
-  3. (3)
-    $b_{i}$ and $b_{j}$ are adjacent if and only if $|i-j|=1\mod n$.
+In fact, a converse to Proposition 6 also holds.
+Proposition 7.
+Suppose $G$ is a (connected) graph. If $\gamma(G\boxempty P_{2})\leq 1$, then $G$ is an outer-cylindrical graph.
+Proof.
+Let $G$ be a (connected) graph such that $\gamma(G\boxempty P_{2})\leq 1$.
+First, if $\gamma(G\boxempty P_{2})=0$, then the classification of planar Cartesian products of graphs given by Behzad and Mahmoodian BeMa1969 implies that $G$ is outerplanar, and so, $G$ is outer-cylindrical.
+Moving forward, we assume $\gamma(G\boxempty P_{2})=1$, which implies that $G$ is not outerplanar.
+Furthermore, there exists a minimal genus embedding of $G\boxempty P_{2}$ on the torus, $\mathbb{T}$, which must be a $2$-cell embedding.
+We work with this fixed embedding for the remainder of this proof.
+Let $G^{+}$ and $G^{-}$ be the two $G$-fiber subgraphs of $G\boxempty P_{2}$.
+Consider the induced embedding of $G^{+}$ on $\mathbb{T}$ and the collection of regions $\mathbb{T}\setminus G^{+}=\{R_{1},\ldots,R_{k}\}$.
+There are two possible cases to consider depending on the topology of these regions.
+Case 1: Suppose each region $R_{i}$, for $1\leq i\leq k$, is a $2$-cell region.
+Since $G^{-}$ is connected, it must be embedded in the interior of some $2$-cell region $R_{i}$.
+Since $G^{+}$ is not outerplanar, there exists some vertex $v^{+}\in V(G^{+})$ that is not on the boundary walk $W$ of the region $R_{i}$.
+However, there exists an edge $e=v^{+}v^{-}$ in a $P_{2}$ fiber connecting $v^{+}$ to the corresponding vertex $v^{-}\in V(G^{-})$, and clearly, $W$ provides an obstruction for $e$ to exist in such an embedding.
+This gives a contradiction.
+Case 2: Suppose there exists some region, say $R_{j}$, that is a non-$2$-cell region.
+First, $R_{j}$ must be the unique non-$2$-cell region in this collection, since otherwise, $G$ would not be connected.
+Then either (i) $R_{j}$ is homeomorphic to a once-punctured torus or (ii) $R_{j}$ is homeomorphic to an open annulus.
+Sub-case (i) implies that $G^{-}$ must be embedded in $R_{j}$, since otherwise, $\gamma(G\boxempty P_{2})=0$.
+This also implies that $G^{+}$ is embedded on a closed disk in $\mathbb{T}$ whose boundary is some walk in $G^{+}$.
+However, we can now apply the same argument used in Case 1 to see that there is a vertex on the interior of this closed disk that must connect via an edge to a vertex in the complement of this closed disk, but such an edge would not exist in this embedding, giving a contradiction for this sub-case.
+So, suppose $R_{j}$ is homeomorphic to an annulus.
+Then $G^{-}$ must be embedded in $R_{j}$, since otherwise, $\gamma(G\boxempty P_{2})=0$.
+Let $W_{1}$ and $W_{2}$ be the two boundary closed walks of $R_{j}$.
+Similar to the argument in Case 1, if there exists $v^{+}\in V(G^{+})$ that is not on $W_{1}\cup W_{2}$, then $W_{1}\cup W_{2}$ forbids the existence of the edge $e=v^{+}v^{-}$, giving a contradiction.
+Thus every vertex of $G^{+}$ must lie on $W_{1}\cup W_{2}$, and so, $G$ is outer-cylindrical, as needed.
+∎
+Collecting our major results provides a proof of our classification theorem that was stated in the introduction.
+Proof of Theorem 1.1.
+Let $G$ be a $3$-connected graph. For the forwards direction, Propositions 5 and 7 collectively give the desired conclusions. For the backwards direction, Proposition 6 gives the desired conclusion. ∎
+To conclude this section, we provide one set of examples of Cartesian products of graphs that are toroidal.
+Example 1.
+For $n\geq 3$, define $C^{2}_{2n}$ as the graph obtained from the $2n$-cycle $C_{2n}$ by adding an edge between every pair of vertices that are a distance two apart in $C_{2n}$.
+See Figure 2 for a standard planar embedding of $C^{2}_{6}$, which also shows this graph is outer-cylindrical.
+For $n\geq 3$, $C^{2}_{2n}$ is known to be $4$-connected; see Ma1980 .
+Thus, $\{C^{2}_{2n}\boxempty P_{2}\}_{n=3}^{\infty}$ provides an infinite class of toroidal Cartesian products of graphs where one factor is $4$-connected.
+Furthermore, by Propositions 4 and 5, we see that $G=C^{2}_{2n}$ realizes the maximal vertex connectivity for toroidal $G\boxempty P_{2}$.
 […]
-Figure 2. The standard planar embedding of $W_{5}$ with vertices labeled.
-Proposition 5.2.
-For $n\geq 3$, each wheel $W_{n}$ is OC. Consequently, there exists an infinite set of OC, $3$-connected graph $\{W_{n}\}_{n=3}^{\infty}$ such that each $W_{n}\boxempty P_{2}$ is toroidal.
-Proof.
-For $n\geq 3$, consider the standard planar embedding of $W_{n}$ as given in Figure 2.
-Then the outer region is a face bound by the $n$-cycle $b_{0}b_{1},\ldots,b_{n-1}$.
-The triangle $cb_{0}b_{1}$ bounds $c$, and so, these two faces contain every vertex, which implies that this embedding is OC.
-The work of Tutte [18] shows that each wheel is simple and $3$-connected.
-By Proposition 5.1, each $W_{n}\boxempty P_{2}$ is toroidal.
-∎
-We can also construct $4$-connected graphs that are OC.
-For $n\geq 4$, define $C^{2}_{n}$ as the graph obtained from the $n$-cycle $C_{n}$ by adding an edge between every pair of vertices that are a distance two apart in $C_{n}$.
-See Figure 3 for a standard planar embedding of $C^{2}_{6}$.
-[…]
-Figure 3. The standard planar embedding of $C^{2}_{6}$ so that the green cycle $v_{1}v_{3}v_{5}$ and red cycle $v_{2}v_{4}v_{6}$ each bound faces for an OC embedding.
-Proposition 5.3.
-For $n\geq 3$, each wheel $C^{2}_{2n}$ is OC.
-Consequently, there exists an infinite set of OC, $4$-connected graph $\{C^{2}_{2n}\}_{n=3}^{\infty}$ such that each $C^{2}_{2n}\boxempty P_{2}$ is toroidal.
-Proof.
-Provide a cyclic labeling of the vertices of cycle $C_{2n}$ as $v_{1},\ldots,v_{2n}$.
-Consider the standard planar embedding of the cycle $C_{2n}$, which partitions the plane into two regions, an inner region and an outer region.
-One can add in the edges of $C^{2}_{2n}$ that are a distance two apart between odd vertices on the inner region and add in the edges of $C^{2}_{2n}$ that are a distance two apart between even vertices on the outer region so that this gives a planar embedding of $C^{2}_{2n}$ where one face (contained inside the inner region) has boundary that contains all the odd vertices and one face (contained in the outer region) has boundary that contains all even vertices; see Figure 3.
-Thus, $C^{2}_{2n}$ is OC.
-For $n\geq 3$, $C^{2}_{2n}$ is known to be $4$-connected; see [9].
-By Proposition 5.1, each $C^{2}_{2n}\boxempty P_{2}$ is toroidal for $n\geq 3$.
-∎
-The work accomplished over the last two sections leads us to the following conjecture.
-Conjecture 5.4.
-Suppose $G$ is $3$-connected. Then $\gamma(G\boxempty H)\leq 1$ if and only if $G$ is OC and $H=P_{2}$.
-We have already proved the backwards direction of Conjecture 5.4.
-We also have shown that if $G$ is $3$-connected and $\gamma(G\boxempty H)\leq 1$, then $H=P_{2}$.
-Furthermore, if a graph $G$ is non-planar, then $\gamma(G\boxempty H)>1$ since $G\boxempty H$ contains two disjoint copies of $G$.
-This leads to the following question:
-Question 5.5.
-If $G$ is a planar, $3$-connected, NOC graph, then must $\gamma(G\boxempty P_{2})>1?$
-In the following section, we provide evidence showing that we don’t expect such $G\boxempty P_{2}$ to be toroidal, further supporting Conjecture 5.4.
-6. $G\boxempty P_{2}$ where $G$ is planar, $3$-connected, and NOC.
-Here, we provide an explicit construction for an infinite set of planar, $3$-connected, NOC graphs $\{H_{n}\}$, and then show that $H_{n}\boxempty P_{2}$ is non-toroidal.
-The construction of $H_{5}$ from a sequence of two vertex splits and one edge addition from the wheel $W_{5}$ is given in Figure 4.
-This process can be generalized to construct $H_{n}$ from $W_{n}$, for $n\geq 5$.
-Moving forward, we refer to the operations of vertex splitting and edge addition collectively as Tutte moves since Tutte used these operations in [18] to build any $3$-connected graph from some wheel graph.
-[…]
-Figure 4. A sequence of three Tutte moves on $W_{5}$ to create a planar, 33-connected NOC graph, $H=H_{5}$.
-Theorem 6.1.
-For each $n\geq 5$, there exist a sequence of three Tutte moves on $W_{n}$ that results in a planar, 3-connected, NOC graph $H_{n}$.
-Proof.
-Consider $W_{n}$, with the vertex labeling for a wheel graph given in Section 5.
-Perform a vertex split on $c$, creating two new vertices $c_{1}$ and $c_{2}$ with an edge connecting them.
-Further, assume $c_{1}$ is degree $3$ and adjacent to $c_{2}$, $b_{0}$, and $b_{1}$, while $c_{2}$ is degree n1n-1 and adjacent to $c_{1}$ and $b_{2},\ldots,b_{n-1}$.
-Next, add an edge ($e$ depicted in red in in part 3 of Figure 4) from $c_{2}$ to $b_{1}$.
-Finally, split $c_{2}$ into $c_{3}$ and $c_{4}$, such that $c_{3}$ is degree $4$ and adjacent to $b_{1}$, $c_{1}$, $b_{n-1}$, and $c_{4}$, while $c_{4}$ is degree n2n-2 and adjacent to $c_{3}$ and $b_{2},\ldots,b_{n-2}$.
-We now check that each $H_{n}$ with $n\geq 5$ has the desired properties.
-Since $W_{n}$ is a $3$-connected graph and Tutte moves preserve being $3$-connected, we also have that $H_{n}$ is a $3$-connected graph.
-In addition, one can start with the standard wheel diagram for $W_{n}$, which is planar, and perform the necessary Tutte moves in a way that preserves planarity to create a planar diagram of $H_{n}$, as shown in Figure 4.
-Next, we see that $H_{5}$ contains a minor that is a minor-minimal NOC graph, specifically $\gamma_{2}$ of Figure 12 in [2].
-This shows $H_{5}$ is NOC.
-Furthermore, we see that $H_{5}$ is a minor of $H_{n}$ (with $n\geq 5$) since one can delete the edges connecting $c_{4}$ to $b_{4},\ldots,b_{n-2}$ from $H_{n}$ to obtain a graph homeomorphic to $H_{5}$.
-Thus, $\gamma_{2}$ is a minor of each such $H_{n}$ with $n\geq 5$.
-Since the class of OC graphs is closed under taking minors, this shows that $H_{n}$ is NOC, for each $n\geq 5$.
-∎
-For the remainder of Section 6, we will refer to the planar, 3-connected, NOC graph depicted in step 4 of Figure 4 as $H=H_{5}$ and $G=\ H\boxempty P_{2}$ to simplify notation.
-Since $H$ is a minor of all graphs obtained via the construction outlined in Theorem 6.1, the genus of $G$ provides a lower bound on the genus of each $H_{n}\boxempty P_{2}$, for $n\geq 5$.
-This motivates us to prove that $G$ is non-toroidal, which we will do by the end of this section in Theorem 6.10.
-First, we note an important corollary of this theorem.
-Corollary 6.2.
-$\{H_{n}\}_{n=5}^{\infty}$ is an infinite set of planar, $3$-connected, NOC graphs such that each $H_{n}\boxempty P_{2}$ is non-toroidal.
-Our corollary also implies that any $3$-connected graph $H^{\prime}$ that contains an $H_{n}$ minor will result in $H^{\prime}\boxempty P_{2}$ being non-toroidal.
-Recall that Tutte [18] showed that every $3$-connected graph is constructed via a finite series of a Tutte moves from some wheel $W_{n}$ for $n\geq 4$.
-Since $H_{n}$ is constructed via just three Tutte moves on $W_{n}$, we see that many $3$-connected graphs will meet the necessary qualities to be a minor of $H$.
-Thus, this work provides further evidence in support of Conjecture 5.4.
-Mirroring the technique used for Theorem 3.4, we now prove the following proposition and lemma that describe the behavior of faces of a $2$-cell embedding of $G$ on the torus.
-Proposition 6.3.
-Suppose $G$ admits a $2$-cell embedding on a torus. This embedding has $|F|=18$ with the following possibilities:
-  1. (1)
-    $f_{3}=6$, $f_{4}=11$, and $f_{6}=1$
-  2. (2)
-    $f_{3}=6$, $f_{4}=10$, and $f_{5}=2$
-  3. (3)
-    $f_{3}=5$, $f_{4}=12$, and $f_{5}=1$, or
-  4. (4)
-    $f_{3}=4$ and $f_{4}=14$.
-Proof.
-Note that, $|E(G)|=34$ and $|V(G)|=16$. Thus, if $G$ admits a $2$-cell embedding on a torus, then this embedding must have $|F|=18$.
-We now consider the possibilities for triangular faces in this embedding.
-Since the girth of $G$ is three, any triangle in this embedding must be bound by a $3$-cycle of $G$.
-We know no edge in the $P_{2}$ fibers is contained in a $3$-cycle, and so, every triangular face in $G$ must bound a $3$-cycle in an $H$-fiber.
-This implies $f_{3}\leq 6$ since each $H$-fiber contains only three $3$-cycles.
-If $f_{3}\leq 3$, then we reach a contradiction because of $2|E(G)|=68=\displaystyle\sum_{i=3}^{\infty}if_{i}$ and $|F|=18$.
-Thus, $4\leq f_{3}\leq f_{6}$.
-From here, one can explicitly check the three possible cases for $f_{3}$ lead to face sets (1)-(4) given in the statement of this Proposition.
-The details are very similar to the proof of Proposition 3.2, which we leave to the reader.
-∎
-Lemma 6.4.
-If $G$ 22-cell embeds on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle in $G$.
-Proof.
-This proof follows the same type of argument given in Lemma 6.4. We leave the details to the reader. ∎
-To prove Theorem 6.10, we now introduce specific notation used to describe $G=H\boxempty P_{2}$ for the remainder of this section.
-We denote the two $H$-fibers of $G$ as $H^{+}$ and $H^{-}$.
-We say that edges $e^{+}\in E(H^{+})$ and $e^{-}\in E(H^{-})$ are corresponding if their images are identical under the projection map $\pi:H\boxempty P_{2}\rightarrow H$.
-Let $E_{p}$ be the set of all edges in the $P_{2}$ fibers of $G$.
-Let $Q$ be the set of all 4-cycles in $G$ such that if $c\in Q$, then $c$ contains two distinct edges in $E_{p}$, one edge in $H^{+}$, and the corresponding edge in $H^{-}$.
-In fact, given any $e\in E_{p}$, the only $4$-cycles that contain $e$ are elements of $Q$.
-Note that $Q$ is in bijective correspondence with the edges of $H$.
-Given an embedding of $G$, we say that faces bound by cycles in $Q$ are mixed quadrilaterals.
-Suppose $G$ 2-cell embeds on some surface $S_{g}$ where for any edge $e\in E_{p}$, we see that $e$ is incident to two (mixed) quadrilateral faces.
-Then we let $Q^{*}\subseteq Q$ represent the set of 4-cycles that collectively bound any mixed quadrilaterals in this embedding.
-We set $E^{+}=Q^{*}\cap H^{+}$, and likewise, $E^{-}=Q^{*}\cap H^{-}$.
-Lemma 6.5.
-Suppose $G$ 2-cell embeds on some surface $S_{g}$ where for any edge $\in E_{p}, ee$ is incident to two (mixed) quadrilateral faces.
-Then the union of elements in $E^{+}$ form a disjoint union of cycles in $H^{+}$ that collectively visit every vertex in $H^{+}$.
-Similarly, the union of elements in $E^{-}$ form a disjoint union of cycles in $H^{-}$ that collectively visit every vertex in $H^{-}$.
-Proof.
-Let $v$ be some vertex in $H^{+}$.
-Then $v$ is incident to a unique edge in $E_{p}$ which is contained in exactly two distinct $4$-cycles in $Q^{*}$.
-By definition, each $Q^{*}$ cycle contains a distinct edge in $E^{+}$ also incident to $v$.
-Thus, every vertex in $H^{+}$ is incident to exactly two distinct edges in $E^{+}$, meaning $E^{+}$ creates a collection of one or more disjoint cycles whose union intersects every vertex in $H^{+}$.
-A similar argument can be replicated for $E^{-}$.
-∎
-We will consider two possible cases for a $2$-cell embedding of $G$ on the torus.
-In the first case, we will assume every edge in $E_{p}$ lies on the boundary of a (mixed) quadrilateral face in this embedding.
-By Proposition 6.3 and the fact that no edge in $E_{p}$ lies on a $3$-cycle of $G$, our second case only needs to consider when only a few edges in $E_{p}$ lies on the boundary of pentagon or hexagon and the rest of the edges of $E_{p}$ lie on the boundary of mixed quadrilaterals.
-We refer to such pentagons as mixed pentagons, and similarly we refer to such hexagons as mixed hexagons.
-Collectively, we refer to mixed quadrilaterals, pentagons, and hexagons of a $2$-cell embedding of $G$ as mixed $n$-gons.
-For now, if we assume every edge in $E_{p}$ lies on the boundary of a (mixed) quadrilateral, then Lemma 6.5 shows the union of elements in $^{+}/EE^{-}$ form a collection of pairwise disjoint cycles $\{C_{i}\}$ in $H$ such that $\sum i=8$.
-This allows us to provide more specification to potential toroidal embeddings of $G$ in terms of such cycle collections.
-Note that, $H$ is a simple graph and one can check that $H$ does not contain a pair of disjoint $4$-cycles.
-Thus, any such collection $\{C_{i}\}$ contains either one $8$-cycle or a $5$-cycle and a $3$-cycle.
-Below, we list these possibilities following the vertex labeling given in Figure 5, and refer to these as cycle sets (of $E^{\pm}$).
-8-cycles in $H$:
-  1. (a)
-    ($v_{1}$, $v_{2}$, $v_{3}$, $v_{6}$, $v_{7}$, $v_{4}$, $v_{8}$, $v_{5}$)
-  2. (b)
-    ($v_{1}$, $v_{2}$, $v_{3}$, $v_{6}$, $v_{7}$, $v_{8}$, $v_{4}$, $v_{5}$)
-  3. (c)
-    ($v_{1}$, $v_{2}$, $v_{6}$, $v_{3}$, $v_{4}$, $v_{7}$, $v_{8}$, $v_{5}$)
-  4. (d)
-    ($v_{1}$, $v_{7}$, $v_{6}$, $v_{2}$, $v_{3}$, $v_{4}$, $v_{8}$, $v_{5}$)
-  5. (e)
-    ($v_{1}$, $v_{2}$, $v_{6}$, $v_{3}$, $v_{4}$, $v_{5}$, $v_{8}$, $v_{7}$)
-Pairs of disjoint 3- and 5-cycles in $H$:
-  1. (f)
-    ($v_{2}$, $v_{3}$, $v_{6}$), ($v_{1}$, $v_{7}$, $v_{4}$, $v_{8}$, $v_{5}$)
-  2. (g)
-    ($v_{4}$, $v_{5}$, $v_{8}$), ($v_{1}$, $v_{2}$, $v_{3}$, $v_{6}$, $v_{7}$)
-[…]
-Figure 5. The graph $H$ with vertex labels.
-Lemma 6.6.
-Suppose $G$ 22-cell embeds on the torus where both copies of the edge $(v_{4},v_{8})$ in the $H$ fibers each lie on the boundary of some mixed $n$-gon face (not necessarily the same one).
-Then $f_{3}=4$ for this embedding and in both of the $H$ fibers, the cycle $(v_{4},v_{7},v_{8})$ or the cycle $(v_{4},v_{5},v_{8})$ can border a triangular face, but not both.
-Proof.
-For sake of contradiction, suppose there exists an embedding of $G$ on the torus with at least five triangles, and both copies of the edge $(v_{4},v_{8})$ in the $H$ fibers each lie on the boundary of some mixed $n$-gon face (not necessarily the same one).
-Note, $G$ contains exactly six $3$-cycles, three in $H^{+}$ and $H^{-}$, respectively.
-Without loss of generality, assume three triangles are bordered by the three $3$-cycles in $H^{+}$: $(v_{2},v_{3},v_{6}),(v_{4},v_{7},v_{8}),$ and $(v_{4},v_{5},v_{8})$.
-Therefore, two triangular faces are incident to edge $v_{4}v_{8}$ in $H^{+}$.
-However, a mixed $n$-gon is also incident to edge $v_{4}v_{8}$ in $H^{+}$.
-Thus, three faces are incident to edge $v_{4}v_{8}$ in $H^{+}$, which contradicts Fact 2.
-The second claim immediately follows from our argument.
-∎
-Proposition 6.7.
-It is not possible for $G$ to $2$-cell embed on the torus where every mixed $n$-gon is a mixed quadrilateral.
-Proof.
-In this proof, we use the term face set in reference to the four possibilities given in Proposition 6.3.
-For sake of contradiction, suppose $G$ does $2$-cell embed on the torus where every mixed $n$-gon is a mixed quadrilateral.
-Then every $e\in E_{p}$ is incident to two mixed quadrilateral faces in this embedding.
-First, note the cycle sets $(a),(b),(d),(f),\text{ and }(g)$ include edge $v_{4}v_{8}$.
-Thus, embeddings corresponding to these cycle sets cannot have face sets 1, 2, and 3 from Proposition 6.3, since otherwise, we contradict Lemma 6.6.
-We now consider cycle sets (a)(g)(a)-(g) as separate cases, keeping in mind the face set restrictions already placed on some of these scenarios.
-Case 1: Consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(a)$.
-As noted in the paragraph above, we only need to consider face set $4$ here, where $f_{3}=4$ and $f_{4}=14$.
-To avoid a contradiction with Lemma 6.6, a triangular face must border the cycle $(v_{2},v_{3},v_{6})$ in each $H$-fiber.
-There must also be a quadrilateral incident to edge $v_{3}v_{4}$ that borders a $4$-cycle in that $H$-fiber.
-However, this face must be adjacent to $v_{3}v_{6}$, and so, $v_{3}v_{6}$ lies on the border to three faces, which contradicts Fact 2.
-The exact same argument can be applied to cycle sets $(b),(f),$ and $(g)$.
-Case 2: Next, consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(c)$.
-Assume a triangle is placed in the face bounded by the 3-cycle ($v_{2},v_{3},v_{6}$) in one of the $H$-fibers.
-There must also be a face incident to edge $v_{6}v_{7}$ that is contained in that $H$-fiber.
-However, this face must be adjacent to either $v_{6}v_{2}$ or $v_{6}v_{3}$, meaning $v_{6}v_{3}$ or $v_{6}v_{2}$ are incident to three faces, which is a contradiction.
-This implies we can have a maximum of two triangular faces per $H$ fiber, and so, face sets 1, 2, and 3 are not possible here.
-Thus, we can assume $f_{3}=4$ and $f_{4}=14$, with triangles bounded by $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in both $H$-fibers.
-Since all remaining faces must be quadrilaterals, edge $v_{4}v_{5}$ must be on the boundary of some face bound by a $4$-cycle in an $H$-fiber.
-However, this means edge $v_{4}v_{7}$ must be incident to three faces, which contradicts Fact 2.
-Case 3: Next, consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(d)$.
-By Lemma 6.6, we see that $f_{3}=4$, $f_{4}=14$, and a triangle must be bound by the cycle $(v_{2},v_{3},v_{6})$ in each of the $H$-fibers.
-Then there must be a quadrilateral in each $H$-fiber that contains edge $v_{2}v_{1}$ on its border.
-However, this face must also contain edge $v_{2}v_{6}$ on its border, and so, $v_{2}v_{6}$ borders three faces, which contradicts Fact 2.
-Case 4: Next, consider an embedding of $G$ where $E^{\pm}$ form the 8-cycle in $(e)$.
-Following the same argument given in Case 3, we see that only face set $4$ is possible where $f_{3}=4$, $f_{4}=14$, and triangles are bound by $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in both $H$-fibers.
-Thus, there must be a quadrilateral whose boundary is a $4$-cycle in an $H$-fiber that contains edge $v_{5}v_{1}$.
-However, this face must be adjacent to either $v_{5}v_{8}$ or $v_{5}v_{4}$, which implies that either $v_{5}v_{8}$ or $v_{5}v_{4}$ is adjacent to three faces, which contradicts Fact 2.
-∎
-The previous case only considers potential toroidal embeddings of $G$ where every mixed $n$-gon is a mixed quadrilateral.
-We now consider potential toroidal embeddings of $G$ where some mixed $n$-gon is not a mixed quadrilateral, and so, either a mixed pentagon or a mixed hexagon must appear in such embeddings.
-Up to exchanging the roles of $H^{\pm}$, any mixed pentagon $p$ must be bound by a $5$-cycle in $G$ that contains two adjacent edges in one $H$-fiber, two edges in $E_{p}$, and one edge in the other $H$-fiber.
-There are two types of mixed hexagons that can occur.
-Up to exchanging roles of $H^{\pm}$, one type of mixed hexagon, $h_{1}$, contains three adjacent edges in $H^{+}$, two edges in $E_{p}$, and one edge in $H^{-}$.
-In the other case, the mixed hexagon $h_{2}$ has two edges in $H^{+}$, two edges in $E_{p}$, and two edges in $H^{-}$.
-Lemma 6.8.
-It is not possible for $G$ to 2-cell embed on the torus with the following conditions:
-  1. (1)
-    Three distinct triangular faces are bound by the 3-cycles in one $H$ fiber.
-  2. (2)
-    Two distinct quadrilateral faces are bound by 4-cycles in the same $H$ fiber.
-Proof.
-For sake of contradiction, suppose we have an embedding of $G$ that meets conditions (1) and (2).
-Without loss of generality, assume conditions (1) and (2) are applied to $H^{+}$.
-Note that, there are five distinct $4$-cycles in $H^{+}$: $(v_{3},v_{4},v_{7},v_{6})$, $(v_{1},v_{2},v_{6},v_{7})$, $(v_{1},v_{7},v_{8},v_{5})$, $(v_{4},v_{5},v_{8},v_{7})$, and $(v_{1},v_{7},v_{4},v_{5})$.
-Furthermore, note that our embedding must admit eight mixed $n$-gons since there are eight edges in $E_{p}$, edges in $E_{p}$ are only bordered by mixed $n$-gons in such an embedding, and each mixed $n$-gon borders exactly two edges in $E_{p}$.
-Consider an embedding with a quadrilateral bound by the 4-cycle $(v_{1},v_{7},v_{8},v_{5})$ in $H^{+}$.
-A mixed $n$-gon must be incident to the edge $e_{p}$ in $E_{p}$ whose vertices correspond with $v_{8}$ in both $H$-fibers.
-Then $e_{p}$ must also be incident to some edge, $e$, in $H^{+}$ incident to $v_{8}$.
-Since triangular faces are bound by the cycles $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in $H^{+}$, we see that $e$ is incident to three faces, which contradicts Fact 2.
-The same type of argument can be applied to an embedding with a quadrilateral bound by the $4$-cycle $(v_{4},v_{5},v_{8},v_{7})$.
-Next, consider an embedding with a quadrilateral bound by the 4-cycle $(v_{1},v_{7},v_{4},v_{5})$ in $H^{+}$.
-Since triangular faces are placed incident to the cycles $(v_{4},v_{7},v_{8})$ and $(v_{4},v_{5},v_{8})$ in $H^{+}$, edges $v_{4}v_{5}$, $v_{4}v_{7}$, and $v_{4}v_{8}$ are incident to exactly two faces just from these two triangles and this one quadrilateral.
-However, there must be a mixed $n$-gon incident to the edge in $E_{p}$ whose endpoints correspond to $v_{4}$ in each $H$-fiber.
-The previous statement forces this mixed $n$-gon to contain the edge $v_{4}v_{3}$ on its boundary.
-Further, this mixed $n$-gon cannot contain the edges $v_{4}v_{5}$, $v_{4}v_{7}$, and $v_{4}v_{8}$ in $H^{+}$ on its boundary, since otherwise, Fact 2 is violated.
-However, any such mixed $n$-gon with these properties will result in a local neighborhood around the $v_{4}$ vertex in $H^{+}$ of this embedding that is not homeomorphic to a disk, which contradicts Fact 1.
-Thus, the two quadrilaterals that bound $4$-cycles in $H^{+}$ must be bound by $(v_{1},v_{2},v_{6},v_{7})$ and $(v_{3},v_{6},v_{7},v_{4})$, respectively.
-Consider the edge $e_{p}\in E_{p}$ whose endpoints correspond to $v_{6}$ in each $H$-fiber.
-This embedding must have a mixed $n$-gon that contains $e_{p}$ on its boundary along with an edge in $H^{+}$ incident to $v_{6}$.
-However, this implies that some edge incident $v_{6}$ is incident to three faces, which contradicts Fact 2.
-∎
-Proposition 6.9.
-It is not possible for $G$ to $2$-cell embed on the torus where some mixed $n$-gon is a mixed pentagon or a mixed hexagon.
-Proof.
-For sake of contradiction, suppose that $G$ does $2$-cell embeds on the torus where some mixed $n$-gon is a mixed pentagon or a mixed hexagon.
-By Proposition 6.3, we see that only face sets (1), (2), or (3) could be used in such an embedding.
-As noted in the proof of Lemma 6.8, our embedding must admit eight mixed $n$-gons.
-Case 1: First, suppose face set (1) was used for this embedding, and so, $f_{3}=6$, $f_{4}=11$, and $f_{6}=1$.
-Thus, there exists one mixed hexagon and seven mixed quadrilaterals in this embedding.
-Since $f_{3}=6$, all three $3$-cycles in both $H$-fibers are bound by triangles.
-Thus, the remaining four quadrilaterals in this embedding must each bound a $4$-cycle in one of the $H$-fibers.
-This implies that at least two quadrilateral faces are bound by $4$-cycles in one of the $H$-fibers.
-Thus, Lemma 6.8 implies that such an embedding is not possible.
-Case 2: Now, suppose face set (2) was used for this embedding, and so, $f_{3}=6$, $f_{4}=10$, and $f_{5}=2$.
-First, all three $3$-cycles in each $H$-fiber must be bound by triangles.
-Such an embedding either has one mixed pentagon and seven mixed quadrilaterals or two mixed pentagons and six mixed quadrilaterals.
-Either way, there must exist at least three quadrilateral faces whose boundaries are $4$-cycles contained in the $H$-fibers.
-Thus, there must exist at least two quadrilateral faces bound by $4$-cycles in either $H^{+}$ or $H^{-}$.
-As a result, Lemma 6.8 shows that such an embedding is not possible.
-Case 3: Finally, suppose face set (3) was used for this embedding, and so, $f_{3}=5$, $f_{4}=12$, and $f_{5}=1$.
-Thus, in one of the $H$-fibers, say $H^{+}$, all three $3$-cycles are bound by triangles.
-In $H^{-}$, exactly two of the three $3$-cycles are bound by triangles.
-We also must have eight mixed $n$-gons, seven of which are mixed quadrilaterals and then one mixed pentagon, $p$.
-This implies that there are five quadrilateral faces whose boundaries are $4$-cycles, where each $4$-cycle is contained in a single $H$-fiber.
-To make sure each of the $13$ edges in each $H$-fiber borders exactly two faces, we must have two quadrilaterals bound by $4$-cycles in $H^{+}$, three quadrilaterals bound by $4$-cycles in $H^{-}$, two edges of $p$ contained in $H^{+}$, and one edge of $p$ contained in $H^{-}$.
-The conditions on $H^{+}$ combined with Lemma 6.8 show that such an embedding is not possible, completing the proof.
-∎
-We can now justify that $G$ has genus two.
-Theorem 6.10.
-Let $G=H\boxempty P_{2}$. Then $\gamma(G)=2$.
-Proof.
-Suppose $G$ does $2$-cell embed on the torus.
-Then by Lemma 6.4, every $n$-gon is bound by an $n$-cycle in $G$.
-By Proposition 6.3, we can only have faces that are triangles, quadrilaterals, pentagons, or hexagons in any such embedding.
-Furthermore, any $e\in E_{p}$ does not lie on a $3$-cycle, and so, any such $e\in E_{p}$ must lie on the boundary of a quadrilateral, pentagon, or hexagon in this embedding, i.e., there must exist mixed $n$-gons in this embedding.
-However, Proposition 6.7 and Proposition 6.9, show that these scenarios are not possible.
-Thus, $\gamma(G)>1$.
-To see that $\gamma(G)=2$, we explicitly construct an embedding of $G$ on $S_{2}$.
-Embed each $H$-fiber of $G=H\boxempty P_{2}$ on a $2$-sphere, following the embedding in Figure 5.
-For each embedding of $H$ on the $2$-sphere, cut out the interiors of three faces, specifically, the ones bound by cycles $(v_{1},v_{2},v_{3},v_{4},v_{5})$, $(v_{3},v_{4},v_{7},v_{6})$, and $(v_{4},v_{5},v_{8})$, respectively.
-Then connect corresponding open patches in each thrice punctured sphere by a tube.
-One can easily embed all the necessary $P_{2}$ fiber edges on these tubes, providing an embedding of $G$ on a surface of genus $2$, as needed.
-∎
-7. Conclusion
-In light of Conjecture 5.4 and the work accomplished here, it remains to prove (or disprove) the following: if $G$ is $3$-connected and $G\boxempty P_{2}$ is toroidal, then $G$ must be OC.
-One might be able to tackle this question by using the list of $38$ minor-minimal NOC graphs given in [2].
-More specifically, one would need an efficient method to check if $G\boxempty P_{2}$ is toroidal where $G$ is one these $38$ minor-minimal NOC graphs.
+Figure 2: The standard planar embedding of $C^{2}_{6}$ so that the green cycle $v_{1}v_{3}v_{5}$ and red cycle $v_{2}v_{4}v_{6}$ each bound faces for an outer-cylindrical embedding.
+Acknowledgements.
+We would like to thank Doug Rall and Thomas Mattman for feedback and suggestions related to this article.
+Funding
+This work was financially supported by the Furman University Department of Mathematics via the Summer Mathematics Undergraduate Research Fellowships.
+Availability of data and materials
+Not applicable.
+Conflict of Interests
+The authors have no relevant financial or non-financial interests to disclose.